### v2

# Hierarchical Vision-Language Prompting for Long-form Video Understanding

Abstract
Long-form videos that span across wide temporal intervals are highly information-redundant and contain multiple distinct events or entities that are often loosely-related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore the use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection and sequence-aware captioning, that can significantly reduce these redundancies. We propose two novel approaches that improve each of aspects, namely Hierarchical Keyframe Selector and Sequential Visual LLM. Our resulting framework termed LVNet achieves state-of-the-art performance across three benchmark LVQA datasets. Our code will be released publicly.
1 Introduction
Video understanding is a long-standing vision problem [Aggarwal2011HumanAA] with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos [allen1994actions, ivanov2000recognition, shi2004propagation, hongeng2004video, Ryoo2006RecognitionOC]. Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos [dataset_tapaswi2016movieqa, videoqaaaai2017, xu2017msvdqavideo].
The recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures [liu2023llava, li2023blip, dai2023instructblip], has resulted in significant improvements across visual question answering (VQA) tasks. These models demonstrate exceptional performance within the image domain, and their video variants [Yu2023SelfChainedIM, papalampidi2023simple, Maaz2023VideoChatGPT] perform similarly on shorter videos, yet demonstrating limited performance on long-form video benchmarks [Mangalam2023EgoSchemaAD, Kahatapitiya2024langrepo, rana2024mvu, rawal2024cinepile]. This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning. An alternate line of works [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, Wang2024VideoAgentLV] adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain. However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames. Such a design choice leading to high compute expense, is questioned in [Buch2022RevisitingT, rana2024mvu, Wang2024VideoAgentLV], and is the key motivation for our exploration of key frame selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair. Moreover, existing captioning methods treats each frame independently while overlooking their sequential nature, which can result in significant redundancies and lack of compatibility between generated captions. This leads to our secondary motivation of exploring sequence-aware captioning strategies for videos.
Figure 1: (left) Our proposed LVNet contains a novel Hierarchical Keyframe Selector module and Sequential Visual LLM strategy followed by an LLM for answer generation. (right) LVNet achieves competitive performance using a fraction of captioned frames highlighting the value of optimal key frame selection.
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) and Sequential Visual LLM (SVL) that performs efficient key-frame selection followed by sequence-aware captioning of individual frames as illustrated in Figure 1. Aligned with prior work [zhang2023llovi, wang2024lifelongmemory, Wang2024VideoAgentLV], the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. The scope of this work focuses on optimizing the prior two stages.
We summarize our key contributions as follows:
  • Long-form Video Understanding framework requiring zero video-level training,
  • Novel Hierarchical Keyframe Selector module,
  • Sequence-aware frame captioning strategy (Sequential Visual LLM module).
Our proposed LVNet achieves state-of-the-art results on three long-form video question answering benchmarks, EgoSchema [Mangalam2023EgoSchemaAD], Next-QA [xiao2021nextqa], and IntentQA [li2023intentqa_cavir], demonstrating strong performance and general applicability of our approach.
2 Related Work
Video Question Answering: Visual question answering (VQA), a relatively recent task, involves generating free-form and open-ended textual content conditioned on an image and natural language text [Agrawal2015VQAVQ]. Its video variant, Video-VQA [Yu2019ActivityNetQAAD] replaces images with videos. While multiple early datasets focus on querying objects or events based on referential and spatial relations [xu2017msvdqavideo, videoqaaaai2017, Yu2019ActivityNetQAAD], later tasks require explicit temporal understanding of sequential events [dataset_lei2018tvqa, lei-etal-2020-tvqaplus, dataset_anetqa]. More recent datasets focus on longer videos containing multiple actions and scenes spread over a wide time interval (termed long-form videos) [dataset_xiao2021nextqa, dataset_li2022causalvidqa] with questions additionally requiring causal reasoning to correctly answer. Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning [dataset_xiao2021nextqa] over long temporal windows [Mangalam2023EgoSchemaAD]. Some works tackling such casual video VQA tasks leverage graph networks to model cross object / event relations [lmappce1_hosseini-etal-2022-knowledge, model_xiao2021hgqa, model_xiao2022vgt]. A more recent line of works integrate large-language models (LLMs) to tackle this task [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, Wang2024VideoAgentLV, rana2024mvu, wang2024lifelongmemory] utilizing the strong reasoning skills of the LLMs. A common aspect is the use of a vision langauge model (VLM) to convert the frame level visual information into natural language. This is in turn input to the LLM which makes a final prediction. In contrast to these methods, our LVNet contains a) a unique Hierarchical Keyframe Selector and b) a novel Sequential Visual LLM strategy for frame captioning. We also highlight that unlike early supervised methods, our LVNet requires no video level training.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video [davis1997mei, Zhao_2017_ICCV] with more recent works focused directly on long-form video question answering [Buch2022RevisitingT, rana2024mvu]. Most similar to our work is [Wang2024VideoAgentLV] which employs an LLM based strategy for video frame selection. However, our proposed Hierarchical Keyframe Selector module differs with unique visual only operations as well as cross-modal operations handling much longer frame sequences extending up to 900 frames.
3 Method
In this section, we present our training-free (i.e. zero-shot) framework for long-form video QA, LVNet. Videos are a dense form of data with even a few seconds long clip being composed of 100s of frames (individual images). In the case of long-form videos, this frame count is even greater. However, the information necessary to answer a given question is often contained in a handful of those frames. Our framework tackles this challenge of selecting an optimal and minimal set of informative frames. We refer to this as keyframe selection. Given such a set of useful frames, we also establish optimal strategies for extracting their information using modern large language models (LLMs), taking into account their sequential nature.
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Sequential Visual LLM (SVL), and an LLM as illustrated in Figure 1. The proposed HKS, an efficient, hierarchical keyframe selector is our key contribution. HKS processes up to 900 frames per video conditioned on a given natural language query in order to select a subset of frames most useful for answering that question. The selected frames are described in natural language using a more expensive, powerful vision language mode (VLM). Our second contribution is a strategy that accounts for the sequential nature of these frames when captioning them using a VLM. Finally, an LLM processes all language descriptions of the selected frames to answer a given query.
3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, rana2024mvu, Wang2024VideoAgentLV] can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the first two stages with emphasis on frame selection.
Under such a framework, our proposed HKS and SVL modules can serve as plug-in modules to replace the frame selection and VLM based frame captioning stages respectively. Our final LLM based answer selection stage operates similar to these prior works.
3.2 Architecture
Consider a long video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, and width respectively. Its paired natural language query is referred as $\mathbf{q}$. Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$. Our goal is to output the suitable response to this query based on information contained in the video. We refer this correct response as $\mathbf{r}$.
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module first processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames are then processed by our SVL stage to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
Our unique contributions here are the Hierarchical Keyframe Selector (HKS) and the Sequential Visual LLM (SVL). We also identify optimal prompts and description templates for the LLM when generating the final response which is also important as suggested by our experiments. The overall architecture is given in Figure 1 (left).
3.3 Hierarchical Keyframe Selector
Figure 2: Detailed architecture of Hierarchical Keyframe Selector: Input frames first go through a Temporal Scene Clustering (TSC) mechanism, which divides the long-video into scenes, enabling per-scene subsampling. Next, a Coarse Keyframe Detector (CKD) is applied to select frames best-aligned with keywords relevant to the query. Finally, a Fine Keyframe detector further subsamples frames, by refining keyword alignements through a templated visual prompting.
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure 2, our HKS comprises of three submodules, namely 1) Temporal Scene Clustering (TSC), 2) Coarse Keyframe Detector (CKD), and 3) Fine Keyframe Detector (FKD). The inputs to HKS contain $T$ frames with each of the three stages TSC, CKD, and FKD reducing this frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling. In contrast, TSC extracts per-frame visual features using a deep neural network followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $k$ frames obtaining a total of $T_{a}=k\times n$. Our iterative clustering procedure is outlined in LABEL:alg:tsc. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based grouping operations. Further details in LABEL:app:icp.
Our reasoning for TSC is that videos, especially long-form ones, contain extensive information redundancy across its temporal dimension. For example, in a long sub-sequence of mostly static frames from a single scene, beyond its first frame the rest may contain little additional information. We hypothesize that our TSC module will utilize frame level feature similarity to cluster the frames, such that each cluster corresponds to a common scene or motion. A single cluster could contain just one frame or significantly more based on frame feature similarities. This leads to a non-uniform sampling of frames across the entire video. In contrast to uniform sampling across the entire videos where each timestamp is treated equally, in our formulation the uniform sampling within segments will allow more frames to be sampled from information heavy temporal regions.
Coarse Keyframe Detector (CKD): Unlike the TSC in prior stage, CKD reasons across both visual and language modalities (using the paired textual query) to further sub-sample the $T_{a}$ frames output from the TSC module. The inputs to this CKD stage are the $T_{a}$ frames and the paired query $\mathbf{q}$. The module generates a set of keywords $K_{\text{CKD}}$, selects $T_{b}$ frames conditioned on these keywords, and outputs the selected $T_{b}$ frames with a confidence value and keyword assigned to each frame.
The CKD contains three elements: a keyword generation strategy, a dual-encoder image-text model, and an iterative algorithm for similarity based confidence assignment. The keyword generation strategy utilizes the given query, q alongside a hand-crafted templating operation or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive langauge image pre-training (CLIP) network from [Ranasinghe2022PerceptualGI]. For confidence assignment, we construct an algorithm as outlined in LABEL:alg:kwimg which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
The keyword generation strategy converts the query $\mathbf{q}$ into a set of keywords, $K_{\text{CKD}}$ where the number of keywords $|K_{\text{CKD}}|\in[8,25]$. The dual-encoder model converts the keyword set $K_{\text{CKD}}$ and the frame set (containing $T_{a}$ frames) into two sets of embeddings. The embeddings of each set are comparable to each other using a cosine similarity distance metric given the training formulation of the CLIP model. These two embedding sets are processed by our confidence assignment algorithm in LABEL:alg:kwimg which assigns each of the $T_{a}$ frames a confidence score and a keyword. Note that the same keyword can be assigned to multiple frames. However, given the much higher frame count compared to the keyword count and the cross-frame redundancy, we do not assign multiple keywords to a single frame.
The key role of this module is language conditioned frame selection. For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query. A single query can also contain multiple different concepts and attributes that must be given attention to construct a correct answer: the keyword generation attempts to capture each of these distinct attributes. On the visual modality, a single frame will also encode multiple concepts and attributes. Our design choice for the spatially aware CLIPpy dual-encoder VLM from [Ranasinghe2022PerceptualGI] is motivated by this nature of individual frames. Finally, confidence assignment takes into account these multiple modes of information within each frame and the query and suitably assigns a confidence score to each frame that reflects its relevance to the query. We also highlight how the confidence scores are directly linked to the related keyword (i.e. reason that makes the frame relevant) leading to better interpretability and the ability to perform further keyword based refinement in later stages.
Fine Keyframe Detector (FKD): In the prior CKD stage, cross-modal selection utilizes a dual-encoder VLM that is constrained by the set of keywords provided and performs limited reasoning at frame level. In contrast, FKD comprises of a) a visual templating module to combines multiple frames and b) an autoregressive VLM that can generate open-ended natural language outputs and is capable of more higher level reasoning. The input to this stage is the set of $F_{b}$ frames, with each frame having an assigned confidence score and keyword.
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure 3 (see Visual Templating column). Each of these visual templated images also contain a subset of the keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see LABEL:app:prompt_fkd for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as the output of the FKD stage, overall resulting in $T_{c}$ frames.
The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at at time). This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage. The eight frames with top confidence scores are grouped into the first visual template, followed by the next eight and so forth. This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage VLM. Our intuition is that such a mechanism allows to best utilize the complimentary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
3.4 Sequential Visual LLM
Our secondary contribution is a strategy for frame caption generation taking into account their sequential nature. We refer to this module as Sequential Visual LLM (SVL).
In contract to the standard approach of independent frame captioning, we explore how per-frame caption generation can be conditioned on captions of previous frames. The VLM used to generate frame descriptions is additionally provided with the caption of one previous frame alongside suitable prompting and templating (details in LABEL:app:vlm).
This is motivated by accounting frame sequential order as well as the presence of redundancy across frames. We discover how this minimal modification to the caption generation strategy results in improved captions, both avoiding caption redundancy and focusing on subtle details in videos (see examples in LABEL:app:vlm).
4 Experiments
In this section, we first discuss our experimental setup followed by quantitative evaluations comparing to existing baselines and ablations of our proposed components. We then present qualitative results for our method and outline some limitations of our approach.
4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [Mangalam2023EgoSchemaAD], NExT-QA [dataset_xiao2021nextqa], and IntentQA [li2023intentqa_cavir]. EgoSchema is a very long-form video question-answering dataset derived from Ego4D. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. We evaluate on both its 500 subset and the 5000 fullset.
NExT-QA is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. We evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA [li2023intentqa_cavir] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions based on 4.3k videos, each with 5 answer choices. These questions are further classified as why?, how? or before/after, based on the appearance of the same keywords in the question.
VLM and LLM Model Choices: In line with prior state-of-the-art work [wang2024lifelongmemory, zhang2023llovi, wang2023vamos] we utilize GPT-4 API for the VLM and LLM in the latter two stages of our framework. This allows fairer comparison against these methods. In the case of the HKS module, we use the same VLM / LLM for keywords used in CKD and VLM used in FKD.
4.2 Quantitative and Qualitative Evaluation
Table 1: Results on EgoSchema [Mangalam2023EgoSchemaAD]: We present comparisons with state-of-the-art models on EgoSchema subset (500-videos) and fullset (5000-videos). LVNet shows a strong zero-shot performance using only a fraction of frames compared to most methods.
Model Captions Subset (%) Fullset (%)
finetuned
ImageViT [papalampidi2023simple] 16 40.8 30.9
ShortViViT [papalampidi2023simple] 32 47.9 31.0
MC-ViT-L [balavzevic2024mcvit]   128+ 62.6 44.4
LongViViT [papalampidi2023simple] 256 56.8 33.3
zero-shot
VIOLET[fu2023violet] 5 - 19.9
mPLUG-Owl [ye2023mplug] 5 - 31.1
VideoAgent [Wang2024VideoAgentLV] 8.4 60.2 54.1
MVU [rana2024mvu] 16 60.3 37.6
SeViLA [yu2024sevila] 32 25.7 22.7
LifelongMemory [wang2024lifelongmemory] 39 26.4 -
InternVideo [wang2022internvideo] 90 - 32.1
FrozenBiLM [yang2022frozenblim] 90 - 26.9
LifelongMemory [wang2024lifelongmemory] 90 68.0 62.1
LLoVi [zhang2023llovi] 180 57.6 50.3
LangRepo [Kahatapitiya2024langrepo] 180 66.2 41.2
Vamos [wang2023vamos] - - 48.3
LVNet (ours) 12 66.0 61.1
We first evaluate on the EgoSchema dataset and present our results in Table 1. Our framework outperforms most prior work to create a stronger baseline with a few frame captions. While LifelongMemory [wang2024lifelongmemory] achieves the absolute best zero-shot performance on the subset at 90 captions (68.0%), its performance drops significantly at 39 captions (26.4%). In contrast, LVNet shows a 66.0% accuracy, merely at 12 captions, even outperforming VideoAgent [Wang2024VideoAgentLV] (60.2% at 8.4 captions) and MVU [rana2024mvu] (60.3% at 16 captions). We achieve 61.1% for the fullest, the highest among the models utilizing approximately 12 frames. Here, we generate captions independently without conditioning on previous captions and present them to the LLM in a system-user-prompt style.
Table 2: Results on NExT-QA [xiao2021nextqa]: We compare LVNet against prior methods on NExT-QA validation set, highlighting standard splits: causal, temporal and descriptive. Our method shows state-of-the-art zero-shot performance, even competing with finetuned models.
Model Captions  Causal (%)  Temporal (%)  Descriptive (%)  All (%)
finetuned
HiTeA [ye2023hitea] 8 62.4 58.3 75.6 63.1
SeViTFiD [kim2023sevit] 10 - - - 60.6
BLIP-2 [li2023blip2] 10 70.1 65.2 80.1 70.1
LLama-VQA [ko2023llama-vqa] 10 72.7 69.2 75.8 72.0
CoVGT [xiao2023covgt] 32 58.8 57.4 69.3 60.0
SeViLA [yu2024sevila] 32 74.2 69.4 81.3 73.8
InternVideo [wang2022internvideo] 90 62.5 58.5 75.8 63.2
MC-ViT-L [balavzevic2024mcvit] 128+ - - - 65.0
Vamos [wang2023vamos] - 72.6 69.6 78.0 72.5
zero-shot
VideoAgent [Wang2024VideoAgentLV] 8.2 72.7 64.5 81.1 71.3
MVU [rana2024mvu] 16 55.4 48.1 64.1 55.2
VFC [momeni2023vfc] 32 45.4 51.6 64.1 51.5
SeViLA [yu2024sevila] 32 61.3 61.5 75.6 63.6
InternVideo [wang2022internvideo] 90 43.4 48.0 65.1 49.1
LLoVi [zhang2023llovi] 90 69.5 61.0 75.6 67.7
LangRepo [Kahatapitiya2024langrepo] 90 64.4 51.4 69.1 60.9
ViperGPT [suris2023vipergpt] - - - - 60.0
LVNet (ours) 12 75.0 65.5 81.5 72.9
We next evaluate on the Next-QA dataset and present results in Table 2. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work by a considerable margin. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy in zero-shot, outperforming VideoAgent [Wang2024VideoAgentLV] (71.3% at 8.2 captions) and LLoVi [zhang2023llovi] (67.7% at 90 captions). It is also extremely competitive with finetuned models such as LLama-VQA [ko2023llama-vqa] (72.0% at 10 captions), SeViLA [yu2024sevila] (73.8% at 32 captions) and Vamos [wang2023vamos] (72.5%), while having no video level training.
Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks. Nevertheless, we highlight how our zero-shot and video level training-free framework is competitive with the best supervised approaches on this dataset. This indicates the promise of utilizing pretrained models, especially those equipped with extensive world knowledge and reasoning skills from alternate modality specific learning (i.e. in our cases image domain VLMs and language domain LLMs).
Table 3: Results on IntentQA [li2023intentqa_cavir]: We compare LVNet against prior methods on IntentQA test set, highlighting standard splits: why?, how? and before/after. Our method shows state-of-the-art performance in zero-shot, with a significant margin.
Model Captions Why? (%) How? (%) Before/After (%) All (%)
finetuned
HQGA [xiao2022hqga] 16 48.2 54.3 41.7 47.7
VGT [xiao2022vgt] 32 51.4 56.0 47.6 51.3
BlindGPT [ouyang2022blindgpt] - 52.2 61.3 43.4 51.6
CaVIR [li2023intentqa_cavir] - 58.4 65.5 50.5 57.6
Vamos [wang2023vamos] - 69.5 70.2 65.0 68.5
zero-shot
SeViLA [yu2024sevila] 32 - - - 60.9
LLoVi [zhang2023llovi] 90 68.4 67.4 51.1 64.0
LangRepo [Kahatapitiya2024langrepo] 90 62.8 62.4 47.8 59.1
LVNet (ours) 12 75.2 71.6 60.8 71.1
We finally evaluate on IntentQA [li2023intentqa_cavir] dataset and report these results in Table 3. Our approach achieves state-of-the-art results on this benchmark as well outperforming all prior work, both zero-shot and finetuned. In fact, LVNet achieves a significant boost (+6.5%) over the closest zero-shot baseline LLoVi [zhang2023llovi] (64.0% at 90 captions), showing a 70.5% accuracy with just 12 frame captions. It also extends the state-of-the-art finetuned performance by +2.0%, compared to Vamos [wang2023vamos] (68.5%).
Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [Mangalam2023EgoSchemaAD]. We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 280 frames. Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM). Such coarse keyframes are then ordered according to confidence values, and grouped into 4 groups of 8 frames each. These groups go through visal templating (i.e. simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
4.3 Ablations
Table 4: Ablations on EgoSchema [Mangalam2023EgoSchemaAD]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
Keyframe sampler Seq. captions #captions Acc. (%)
Uniform Sampling (LLoVi) 180 57.6
HKS (w/o SVL) 12 63.0
HKS 12 63.4
Visual Templating order Acc. (%)
Temporal order 62.9
Confidence order 63.4
LLM Acc. (%)
GPT-3.5 59.2
GPT-4 61.1
GPT-4o 63.4
Frames 6.4 8.4 9 11 12 16
VideoAgent 58.4 60.2 - 57.4 - -
LVNet (ours) - - 64.6 - 66.0 65.8
In this section, we present ablations to key design decisions such as #frames, #words, prompting, and hierarchical processing. The results for these ablations are presented in Table 4. In all ablations, we consider a subset of 150 videos from EgoSchema [Mangalam2023EgoSchemaAD] for evaluation.
Keyframe sampler:
We evaluate our Hierarchical Keyframe Sampling strategy (HKS) against the established uniform sampling baseline from [zhang2023llovi]. Our results clearly indicate the importance of our HKS module.
Sequential captioning:
In our SVL module, we consider a sequential captioner, that considers previous captions when captioning the current frame. It outperforms the common parallel frame-wise captioning by $+0.4\%$ showing that the SVL can generate better captions, possibly reducing redundant information among captions and enabling a form of temporal modeling.
Visual Templating order:
In Visual Templating, grouping the frames based on confidence is better than preserving temporal order. It can allow sampling of challenging keyword frames, instead of only selecting high confidence keyframes.
Choice of LLM:
Here, we see that more-recent GPT-4o outperforms both GPT-4 and GPT-3.5. Since GPT-4o is also cheaper compared to GPT-4, we consider it as our default choice.
Number of frame captions:
We ablate different number of frame captions used in our setup compared to VideoAgent [Wang2024VideoAgentLV]. Our method shows more-stable performance.
5 Conclusion
We proposed an approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art results across 3 benchmarks datasets. Our propsoed Hierarchical Keyframe Selector and Sequential Visual LLM modules demonstrate the usefulness of keyframe selection and sequence-aware frame captioning. We also highlight the zero-shot capability for long-form video understanding of our LVNet framework requiring zero video level training, and our experiments demonstrate its clear advantage over prior work.

### v3

@@
-Hierarchical Vision-Language Prompting for Long-form Video Understanding
+Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA
 Abstract
-Long-form videos that span across wide temporal intervals are highly information-redundant and contain multiple distinct events or entities that are often loosely-related.
+Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related.
 Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames.
@@
 Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant.
-Questioning these decision choices, we explore optimal strategies for key-frame selection and sequence-aware captioning, that can significantly reduce these redundancies.
-We propose two novel approaches that improve each of aspects, namely Hierarchical Keyframe Selector and Sequential Visual LLM.
-Our resulting framework termed LVNet achieves state-of-the-art performance across three benchmark LVQA datasets.
-Our code will be released publicly.
+Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector.
+Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, IntentQA, NExT-QA.
+The code can be found at https://github.com/jongwoopark7978/LVNet
 1 Introduction
-Video understanding is a long-standing vision problem [Aggarwal2011HumanAA] with numerous real-world applications.
-It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos [allen1994actions, ivanov2000recognition, shi2004propagation, hongeng2004video, Ryoo2006RecognitionOC].
-Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos [dataset_tapaswi2016movieqa, videoqaaaai2017, xu2017msvdqavideo].
-The recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures [liu2023llava, li2023blip, dai2023instructblip], has resulted in significant improvements across visual question answering (VQA) tasks.
-These models demonstrate exceptional performance within the image domain, and their video variants [Yu2023SelfChainedIM, papalampidi2023simple, Maaz2023VideoChatGPT] perform similarly on shorter videos, yet demonstrating limited performance on long-form video benchmarks [Mangalam2023EgoSchemaAD, Kahatapitiya2024langrepo, rana2024mvu, rawal2024cinepile].
+Video understanding is a long-standing vision problem [1] with numerous real-world applications.
+It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos [3, 15, 33, 13, 32].
+Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos [34, 51, 44].
+The recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures [23, 21, 7], has resulted in significant improvements across visual question answering (VQA) tasks.
+These models demonstrate exceptional performance within the image domain, and their video variants [47, 28, 27] perform similarly on shorter videos, yet demonstrating limited performance on long-form video benchmarks [24, 16, 30, 31].
 This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning.
-An alternate line of works [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, Wang2024VideoAgentLV] adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain.
+An alternate line of works [52, 36, 16, 37] adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain.
 However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames.
-Such a design choice leading to high compute expense, is questioned in [Buch2022RevisitingT, rana2024mvu, Wang2024VideoAgentLV], and is the key motivation for our exploration of key frame selection, i.e.
+Such a design choice leading to high compute expense, is questioned in [4, 30, 37], and is the key motivation for our exploration of key frame selection, i.e.
 identifying a minimal set of frames most useful for correctly answering a given video-question pair.
-Moreover, existing captioning methods treats each frame independently while overlooking their sequential nature, which can result in significant redundancies and lack of compatibility between generated captions.
-This leads to our secondary motivation of exploring sequence-aware captioning strategies for videos.
-Figure 1: (left) Our proposed LVNet contains a novel Hierarchical Keyframe Selector module and Sequential Visual LLM strategy followed by an LLM for answer generation.
-(right) LVNet achieves competitive performance using a fraction of captioned frames highlighting the value of optimal key frame selection.
-Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) and Sequential Visual LLM (SVL) that performs efficient key-frame selection followed by sequence-aware captioning of individual frames as illustrated in Figure 1.
-Aligned with prior work [zhang2023llovi, wang2024lifelongmemory, Wang2024VideoAgentLV], the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair.
+Figure 1: (left) Our proposed LVNet contains a novel Hierarchical Keyframe Selector module to select keyframes and followed by VLM and LLM for caption and answer generation.
+Hierarchical Keyframe Selector initially begins by processing dense frames and keywords with the lighter module and progressively exploits heavier and more performance-oriented modules on a small set of frames, resulting from the reduction of keyframe candidates, to ensure efficient computation.
+(right) LVNet achieves state-of-the-art performance on the EgoSchema subset while utilizing only a fraction of captioned frames.
+In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions, outperforming VideoAgent, a model using a similar-scale captions, by 8%.
+A more detailed analysis of accuracy vs.
+the number of captions is provided in 4.3.
+Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Figure 1.
+Aligned with prior work [52, 39, 37], the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair.
 The scope of this work focuses on optimizing the prior two stages.
 We summarize our key contributions as follows:
-  • Long-form Video Understanding framework requiring zero video-level training,
-  • Novel Hierarchical Keyframe Selector module,
-  • Sequence-aware frame captioning strategy (Sequential Visual LLM module).
-Our proposed LVNet achieves state-of-the-art results on three long-form video question answering benchmarks, EgoSchema [Mangalam2023EgoSchemaAD], Next-QA [xiao2021nextqa], and IntentQA [li2023intentqa_cavir], demonstrating strong performance and general applicability of our approach.
+  • Novel Hierarchical Keyframe Selector module which are composed of three submodules: Temporal Scene Clustering (TSC), Coarse Keyframe Detector (CKD), and Fine Keyframe Detector (FKD).
+    • Temporal Scene Clustering performs non-uniform frame sampling by clustering visually similar frames, minimizing redundancy in long videos and capturing key scenes.
+The lightweight module is used to efficiently filter dense frames.
+    • Coarse Keyframe Detector generates keywords, representing atomic activities, using the given query and LLM.
+It assigns confidence scores to frames based on keyword relevance, sub-sampling high-confidence frames for improved interpretability over visual-only selection.
+    • Fine Keyframe Detector refines frame selection by combining multiple frames using visual templating and Visual Language Model (VLM), enabling higher-level reasoning and natural language output for improved accuracy over the CKD’s keyword-based selection.
+  • Long-form Video Understanding framework requiring zero video-level training.
+Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], IntentQA[22], and NExT-QA[41] as described in 4.2.
+This demonstrates the strong performance and general applicability of our approach.
 2 Related Work
-Video Question Answering: Visual question answering (VQA), a relatively recent task, involves generating free-form and open-ended textual content conditioned on an image and natural language text [Agrawal2015VQAVQ].
-Its video variant, Video-VQA [Yu2019ActivityNetQAAD] replaces images with videos.
-While multiple early datasets focus on querying objects or events based on referential and spatial relations [xu2017msvdqavideo, videoqaaaai2017, Yu2019ActivityNetQAAD], later tasks require explicit temporal understanding of sequential events [dataset_lei2018tvqa, lei-etal-2020-tvqaplus, dataset_anetqa].
-More recent datasets focus on longer videos containing multiple actions and scenes spread over a wide time interval (termed long-form videos) [dataset_xiao2021nextqa, dataset_li2022causalvidqa] with questions additionally requiring causal reasoning to correctly answer.
-Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning [dataset_xiao2021nextqa] over long temporal windows [Mangalam2023EgoSchemaAD].
-Some works tackling such casual video VQA tasks leverage graph networks to model cross object / event relations [lmappce1_hosseini-etal-2022-knowledge, model_xiao2021hgqa, model_xiao2022vgt].
-A more recent line of works integrate large-language models (LLMs) to tackle this task [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, Wang2024VideoAgentLV, rana2024mvu, wang2024lifelongmemory] utilizing the strong reasoning skills of the LLMs.
+Video Question Answering: Visual question answering (VQA), a relatively recent task, involves generating free-form and open-ended textual content conditioned on an image and natural language text [2].
+Its video variant, Video-VQA [49] replaces images with videos.
+While multiple early datasets focus on querying objects or events based on referential and spatial relations [44, 51, 49], later tasks require explicit temporal understanding of sequential events [18, 19, 50].
+More recent datasets focus on longer videos containing multiple actions and scenes spread over a wide time interval (termed long-form videos) [41, 20] with questions additionally requiring causal reasoning to correctly answer.
+Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning [41] over long temporal windows [24].
+Some works tackling such casual video VQA tasks leverage graph networks to model cross object / event relations [14, 42, 43].
+A more recent line of works integrate large-language models (LLMs) to tackle this task [52, 36, 16, 37, 30, 39, 9] utilizing the strong reasoning skills of the LLMs.
 A common aspect is the use of a vision langauge model (VLM) to convert the frame level visual information into natural language.
 This is in turn input to the LLM which makes a final prediction.
-In contrast to these methods, our LVNet contains a) a unique Hierarchical Keyframe Selector and b) a novel Sequential Visual LLM strategy for frame captioning.
-We also highlight that unlike early supervised methods, our LVNet requires no video level training.
-Frame Selection in Videos: The task of frame selection in videos has been long explored in video [davis1997mei, Zhao_2017_ICCV] with more recent works focused directly on long-form video question answering [Buch2022RevisitingT, rana2024mvu].
-Most similar to our work is [Wang2024VideoAgentLV] which employs an LLM based strategy for video frame selection.
+Unlike these methods, our LVNet incorporates a unique Hierarchical Keyframe Selector that progressively reduces the number of keyframe candidates.
+Lighter modules are applied to dense frames, while heavier, more performance-focused modules are applied to a small set of frames for efficient computation.
+Additionally, LVNet does not require video-level training, unlike earlier supervised approaches.
+Frame Selection in Videos: The task of frame selection in videos has been long explored in video [8, 53] with more recent works focused directly on long-form video question answering [4, 30].
+Most similar to our work is [37] which employs an LLM based strategy for video frame selection.
 However, our proposed Hierarchical Keyframe Selector module differs with unique visual only operations as well as cross-modal operations handling much longer frame sequences extending up to 900 frames.
@@
 Given such a set of useful frames, we also establish optimal strategies for extracting their information using modern large language models (LLMs), taking into account their sequential nature.
-Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Sequential Visual LLM (SVL), and an LLM as illustrated in Figure 1.
-The proposed HKS, an efficient, hierarchical keyframe selector is our key contribution.
-HKS processes up to 900 frames per video conditioned on a given natural language query in order to select a subset of frames most useful for answering that question.
-The selected frames are described in natural language using a more expensive, powerful vision language mode (VLM).
-Our second contribution is a strategy that accounts for the sequential nature of these frames when captioning them using a VLM.
-Finally, an LLM processes all language descriptions of the selected frames to answer a given query.
+Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure 1.
+The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work.
+First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes using ResNet-18, a lightweight module that measures visual correspondence between frames.
+Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords using CLIP-B/16.
+Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM.
+Finally, an LLM processes the language descriptions of the selected frames to answer a given query.
 3.1 Background
-Recent approaches utilizing LLMs for long video question answering (LVQA) [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, rana2024mvu, Wang2024VideoAgentLV] can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation.
+Recent approaches utilizing LLMs for long video question answering (LVQA) [52, 36, 16, 30, 37] can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation.
 Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g.
 frame selection in some is simply uniform sampling).
-In our work, we also follow this structure, but we focus on improving the first two stages with emphasis on frame selection.
-Under such a framework, our proposed HKS and SVL modules can serve as plug-in modules to replace the frame selection and VLM based frame captioning stages respectively.
-Our final LLM based answer selection stage operates similar to these prior works.
+In our work, we also follow this structure, but we focus on improving the frame selection stage.
+Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
+Figure 2: Detailed architecture of Hierarchical Keyframe Selector: Input frames first go through a Temporal Scene Clustering (TSC) mechanism, which divides the long-video into scenes, enabling per-scene subsampling.
+Next, a Coarse Keyframe Detector (CKD) is applied to select frames best-aligned with keywords relevant to the query.
+Finally, a Fine Keyframe detector further subsamples frames, by refining keyword alignements through a templated visual prompting.
 3.2 Architecture
@@
 Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$.
-The HKS module first processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$.
-Each of these $T^{\prime}$ frames are then processed by our SVL stage to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$.
+The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$.
+Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$.
 Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
-Our unique contributions here are the Hierarchical Keyframe Selector (HKS) and the Sequential Visual LLM (SVL).
-We also identify optimal prompts and description templates for the LLM when generating the final response which is also important as suggested by our experiments.
+Our unique contributions here are the Hierarchical Keyframe Selector (HKS) and optimal prompts and description templates for the LLM when generating the final response which is also important as suggested by our experiments.
 The overall architecture is given in Figure 1 (left).
 3.3 Hierarchical Keyframe Selector
-Figure 2: Detailed architecture of Hierarchical Keyframe Selector: Input frames first go through a Temporal Scene Clustering (TSC) mechanism, which divides the long-video into scenes, enabling per-scene subsampling.
-Next, a Coarse Keyframe Detector (CKD) is applied to select frames best-aligned with keywords relevant to the query.
-Finally, a Fine Keyframe detector further subsamples frames, by refining keyword alignements through a templated visual prompting.
+Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [24].
+We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 390 frames.
+Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM).
+Such coarse keyframes are then ranked based on the combination of confidence value and temporal span, and grouped into four sets, each containing eight frames.
+These sets are then processed through visual templating (i.e.
+simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
 We now describe our proposed Hierarchical Keyframe Selector (HKS) module.
@@
 In contrast, TSC extracts per-frame visual features using a deep neural network followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets.
-Within each of the n sets, we uniformly sample $k$ frames obtaining a total of $T_{a}=k\times n$.
-Our iterative clustering procedure is outlined in LABEL:alg:tsc.
+Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$.
+Our iterative clustering procedure is outlined in Algorithm 1.
 It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based grouping operations.
-Further details in LABEL:app:icp.
 Our reasoning for TSC is that videos, especially long-form ones, contain extensive information redundancy across its temporal dimension.
@@
 The keyword generation strategy utilizes the given query, q alongside a hand-crafted templating operation or an LLM to select or generate suitable keywords.
-The dual-encoder image-text model uses a spatially aware contrastive langauge image pre-training (CLIP) network from [Ranasinghe2022PerceptualGI].
-For confidence assignment, we construct an algorithm as outlined in LABEL:alg:kwimg which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
+The dual-encoder image-text model uses a spatially aware contrastive langauge image pre-training (CLIP) network from [29].
+For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
 The keyword generation strategy converts the query $\mathbf{q}$ into a set of keywords, $K_{\text{CKD}}$ where the number of keywords $|K_{\text{CKD}}|\in[8,25]$.
@@
 The embeddings of each set are comparable to each other using a cosine similarity distance metric given the training formulation of the CLIP model.
-These two embedding sets are processed by our confidence assignment algorithm in LABEL:alg:kwimg which assigns each of the $T_{a}$ frames a confidence score and a keyword.
+These two embedding sets are processed by our confidence assignment algorithm in Appendix A.2 which assigns each of the $T_{a}$ frames a confidence score and a keyword.
 Note that the same keyword can be assigned to multiple frames.
@@
 On the visual modality, a single frame will also encode multiple concepts and attributes.
-Our design choice for the spatially aware CLIPpy dual-encoder VLM from [Ranasinghe2022PerceptualGI] is motivated by this nature of individual frames.
+Our design choice for the spatially aware CLIPpy dual-encoder VLM from [29] is motivated by this nature of individual frames.
 Finally, confidence assignment takes into account these multiple modes of information within each frame and the query and suitably assigns a confidence score to each frame that reflects its relevance to the query.
 We also highlight how the confidence scores are directly linked to the related keyword (i.e.
-reason that makes the frame relevant) leading to better interpretability and the ability to perform further keyword based refinement in later stages.
+reason that makes the frame relevant), leading to better interpretability and the ability to perform further keyword-based refinement in later stages.
 Fine Keyframe Detector (FKD): In the prior CKD stage, cross-modal selection utilizes a dual-encoder VLM that is constrained by the set of keywords provided and performs limited reasoning at frame level.
-In contrast, FKD comprises of a) a visual templating module to combines multiple frames and b) an autoregressive VLM that can generate open-ended natural language outputs and is capable of more higher level reasoning.
-The input to this stage is the set of $F_{b}$ frames, with each frame having an assigned confidence score and keyword.
+In contrast, FKD uses a visual templating module to combine multiple frames and uses VLM to generate open-ended natural language output through higher-level reasoning.
+The input in this stage is the set of $F_{b}$ frames, with each frame having an assigned confidence score and keyword.
 Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame.
@@
 Each of these visual templated images also contain a subset of the keywords that correspond to their 8 images.
-These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see LABEL:app:prompt_fkd for details) are input to the VLM.
+These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.3 for details) are input to the VLM.
 The output of the VLM is used to select a subset of each 8 image group.
 These frames are collected as the output of the FKD stage, overall resulting in $T_{c}$ frames.
-The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at at time).
-This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage.
+The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at time).
+This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage and timestamps.
 The eight frames with top confidence scores are grouped into the first visual template, followed by the next eight and so forth.
-This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage VLM.
-Our intuition is that such a mechanism allows to best utilize the complimentary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
-3.4 Sequential Visual LLM
-Our secondary contribution is a strategy for frame caption generation taking into account their sequential nature. We refer to this module as Sequential Visual LLM (SVL).
-In contract to the standard approach of independent frame captioning, we explore how per-frame caption generation can be conditioned on captions of previous frames.
-The VLM used to generate frame descriptions is additionally provided with the caption of one previous frame alongside suitable prompting and templating (details in LABEL:app:vlm).
-This is motivated by accounting frame sequential order as well as the presence of redundancy across frames.
-We discover how this minimal modification to the caption generation strategy results in improved captions, both avoiding caption redundancy and focusing on subtle details in videos (see examples in LABEL:app:vlm).
+This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage.
+After that, we temporally reorder some image sets with low confidence scores to cover keyframes distributed across long-range segments, while the sets with high confidence scores concentrate on keyframes in short-range segments.
+A total of 16 low-score frames are temporally reordered in this process.
+The algorithm is described in Algorithm 3 and the prompting technique is explained in Appendix A.3.
+Our intuition is that such a mechanism allows one to best utilize the complementary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
+3.4 Comparison with Other Keyframe Selection Methods:
+We aim to highlight the main advantage of the Hierarchical Keyframe Selector over other existing keyframe selection methods.
+Models like VideoAgent and TraveLER provide useful comparisons, as they work with a similar-scale number of captions.
+VideoAgent and TraveLER rely on uniform frame selection in the first iteration without analyzing the entire video.
+They identify important segments based solely on these initial frames and the LLM’s response, which can be problematic if the initial uniformly selected frames are not representative of the entire video or if the LLM misinterprets the captions and prompts.
+In such cases, the LLM might incorrectly identify segments for further analysis.
+If the LLM fails to pinpoint the correct segment initially, the entire process can break down because subsequent frames will be similar to the first set, leading the LLM to continuously select frames within or near the initial segment.
+Additionally, for videos that are as challenging or more difficult than EgoSchema in terms of temporal complexity and activities, these models may require numerous iterations to finalize keyframes selection.
+This results in higher computational and latency costs, as it necessitates numerous runs of resource-intensive VLM and LLM models.
+In contrast, our method analyzes the entire video with high frame rates using a lightweight ResNet-18 [12] and segments the video non-uniformly based on scene continuity.
+We then select several frames in each segment by measuring feature similarity between frame features and keywords using the CLIP-B/16 (0.12B) [29] which is lighted than VideoAgent’s EVA-CLIP-8Bplus (8B).
+By reviewing the entire video and non-uniformly selecting keyframes based on scene continuity and similarity scores, these keyframes accurately represent the question-based important frames distribution in the entire video.
+Furthermore, we use VLM for a fine-grained selection of keyframes, improving keyframe selection when CLIP-B/16 struggles to understand detailed atomic activities in the frames.
+By hierarchically segmenting the video with different modules, the resulting segments and keyframes are more reliable than those from VideoAgent.
+Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.
 4 Experiments
@@
 Datasets are used purely for evaluation.
-We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [Mangalam2023EgoSchemaAD], NExT-QA [dataset_xiao2021nextqa], and IntentQA [li2023intentqa_cavir].
-EgoSchema is a very long-form video question-answering dataset derived from Ego4D.
+We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], IntentQA [22], and NExT-QA [41].
+The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions.
 For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip.
-We evaluate on both its 500 subset and the 5000 fullset.
-NExT-QA is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer.
-We evaluate on its validation set containing 4996 questions over 570 videos.
-The third dataset, IntentQA [li2023intentqa_cavir] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions.
-It consists of 16k multiple-choice questions based on 4.3k videos, each with 5 answer choices.
-These questions are further classified as why?, how?
-or before/after, based on the appearance of the same keywords in the question.
-VLM and LLM Model Choices: In line with prior state-of-the-art work [wang2024lifelongmemory, zhang2023llovi, wang2023vamos] we utilize GPT-4 API for the VLM and LLM in the latter two stages of our framework.
-This allows fairer comparison against these methods.
-In the case of the HKS module, we use the same VLM / LLM for keywords used in CKD and VLM used in FKD.
-4.2 Quantitative and Qualitative Evaluation
-Table 1: Results on EgoSchema [Mangalam2023EgoSchemaAD]: We present comparisons with state-of-the-art models on EgoSchema subset (500-videos) and fullset (5000-videos).
-LVNet shows a strong zero-shot performance using only a fraction of frames compared to most methods.
-Model Captions Subset (%) Fullset (%)
-finetuned
-ImageViT [papalampidi2023simple] 16 40.8 30.9
-ShortViViT [papalampidi2023simple] 32 47.9 31.0
-MC-ViT-L [balavzevic2024mcvit]   128+ 62.6 44.4
-LongViViT [papalampidi2023simple] 256 56.8 33.3
-zero-shot
-VIOLET[fu2023violet] 5 - 19.9
-mPLUG-Owl [ye2023mplug] 5 - 31.1
-VideoAgent [Wang2024VideoAgentLV] 8.4 60.2 54.1
-MVU [rana2024mvu] 16 60.3 37.6
-SeViLA [yu2024sevila] 32 25.7 22.7
-LifelongMemory [wang2024lifelongmemory] 39 26.4 -
-InternVideo [wang2022internvideo] 90 - 32.1
-FrozenBiLM [yang2022frozenblim] 90 - 26.9
-LifelongMemory [wang2024lifelongmemory] 90 68.0 62.1
-LLoVi [zhang2023llovi] 180 57.6 50.3
-LangRepo [Kahatapitiya2024langrepo] 180 66.2 41.2
-Vamos [wang2023vamos] - - 48.3
-LVNet (ours) 12 66.0 61.1
-We first evaluate on the EgoSchema dataset and present our results in Table 1.
-Our framework outperforms most prior work to create a stronger baseline with a few frame captions.
-While LifelongMemory [wang2024lifelongmemory] achieves the absolute best zero-shot performance on the subset at 90 captions (68.0%), its performance drops significantly at 39 captions (26.4%).
-In contrast, LVNet shows a 66.0% accuracy, merely at 12 captions, even outperforming VideoAgent [Wang2024VideoAgentLV] (60.2% at 8.4 captions) and MVU [rana2024mvu] (60.3% at 16 captions).
-We achieve 61.1% for the fullest, the highest among the models utilizing approximately 12 frames.
-Here, we generate captions independently without conditioning on previous captions and present them to the LLM in a system-user-prompt style.
-Table 2: Results on NExT-QA [xiao2021nextqa]: We compare LVNet against prior methods on NExT-QA validation set, highlighting standard splits: causal, temporal and descriptive.
-Our method shows state-of-the-art zero-shot performance, even competing with finetuned models.
-Model Captions  Causal (%)  Temporal (%)  Descriptive (%)  All (%)
-finetuned
-HiTeA [ye2023hitea] 8 62.4 58.3 75.6 63.1
-SeViTFiD [kim2023sevit] 10 - - - 60.6
-BLIP-2 [li2023blip2] 10 70.1 65.2 80.1 70.1
-LLama-VQA [ko2023llama-vqa] 10 72.7 69.2 75.8 72.0
-CoVGT [xiao2023covgt] 32 58.8 57.4 69.3 60.0
-SeViLA [yu2024sevila] 32 74.2 69.4 81.3 73.8
-InternVideo [wang2022internvideo] 90 62.5 58.5 75.8 63.2
-MC-ViT-L [balavzevic2024mcvit] 128+ - - - 65.0
-Vamos [wang2023vamos] - 72.6 69.6 78.0 72.5
-zero-shot
-VideoAgent [Wang2024VideoAgentLV] 8.2 72.7 64.5 81.1 71.3
-MVU [rana2024mvu] 16 55.4 48.1 64.1 55.2
-VFC [momeni2023vfc] 32 45.4 51.6 64.1 51.5
-SeViLA [yu2024sevila] 32 61.3 61.5 75.6 63.6
-InternVideo [wang2022internvideo] 90 43.4 48.0 65.1 49.1
-LLoVi [zhang2023llovi] 90 69.5 61.0 75.6 67.7
-LangRepo [Kahatapitiya2024langrepo] 90 64.4 51.4 69.1 60.9
-ViperGPT [suris2023vipergpt] - - - - 60.0
-LVNet (ours) 12 75.0 65.5 81.5 72.9
-We next evaluate on the Next-QA dataset and present results in Table 2.
+The second dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions.
+It consists of 16k multiple-choice questions which are classified as Why?, How?
+or Before/After (B./A.).
+The third dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer.
+These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos.
+Model Choices: For the HKS module, we use the ResNet-18 [12] for the TSC, CLIP-B/16 [29] for the CKD and GPT-4o for the FKD.
+We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T.
+This makes them well-suited for filtering dense frames efficiently.
+In line with previous state-of-the-art work [39, 52, 36], we employ GPT API, especially GPT-4o, for both VLM and LLM.
+This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4.
+GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework.
+Table 1: Combined Results on IntentQA [22], EgoSchema [24], and NExT-QA [41].
+We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits.
+The models are ordered based on the number of captions processed before answering the question.
+LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on the three datasets, respectively, using just 12 frames compared to the models using the similar number of captions.
+Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
+Model IntentQA EgoSchema NExT-QA
+Cap. Why? (%) How? (%) B./A. (%) All (%) Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%)
+Vamos [36] - - - - - - 48.3 - - - - -
+IG-VLM [17] - - - - 65.3 - 59.8 - 69.8 63.6 74.7 68.6
+VideoLLaMA 2 [5] - - - - - - 53.3 - - - - -
+InternVideo2 [38] - - - - - - 60.2 - - - - -
+Tarsier [35] - - - - - - 61.7 - - - - 79.2
+VIOLET [10] - - - - - 5 19.9 - - - - -
+mPLUG-Owl [46] - - - - - 5 31.1 - - - - -
+VideoAgent [37] - - - - - 8.4 54.1 8.2 72.7 64.5 81.1 71.3
+MVU [30] - - - - - 16 37.6 16 55.4 48.1 64.1 55.2
+MoReVQA [25] - - - - - 30 51.7 16 70.2 64.6 - 69.2
+TraveLER [37] - - - - - - - $\sim$25 70.0 60.5 78.2 68.2
+VFC [26] - - - - - - - 32 45.4 51.6 64.1 51.5
+SeViLA [48] 32 - - - 60.9 32 22.7 32 61.3 61.5 75.6 63.6
+ProViQ [6] - - - - - 60 57.1 60 - - - 64.6
+VideoTree [40] $\leq$56 - - - 66.9 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5
+FrozenBiLM [45] - - - - - 90 26.9 - - - - -
+LifelongMemory [39] - - - - - 90 62.1 - - - - -
+LangRepo [16] 90 62.8 62.4 47.8 59.1 180 41.2 90 64.4 51.4 69.1 60.9
+LLoVi [52] 90 68.4 67.4 51.1 64.0 180 50.3 90 69.5 61.0 75.6 67.7
+LVNet (ours) 12 75.0 74.4 62.1 71.7 12 61.1 12 75.0 65.5 81.5 72.9
+4.2 Evaluation
+Quantitative Results:
+We evaluate LVNet on the IntentQA, EgoSchema, and NExT-QA dataset and present our results in Table 1.
+Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining.
+Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency.
+For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions.
+This result outperforms VideoAgent, the next best model using 8.4 captions, by 7%.
+To ensure a fair comparison, we ablate the number of captions for LVNet in LABEL:tab:ablation:numberofframecaptions, observing that LVNet with 8 captions still outperforms VideoAgent with 8.4 captions by 4.2% in the EgoSchema subset.
+The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
+We next evaluate on the IntentQA[22] dataset.
+Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames.
+In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
+We finally evaluate on the NExT-QA[41] dataset.
 This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs.
-Our approach achives state-of-the-art performance on this benchmark outperforming prior work by a considerable margin.
-In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy in zero-shot, outperforming VideoAgent [Wang2024VideoAgentLV] (71.3% at 8.2 captions) and LLoVi [zhang2023llovi] (67.7% at 90 captions).
-It is also extremely competitive with finetuned models such as LLama-VQA [ko2023llama-vqa] (72.0% at 10 captions), SeViLA [yu2024sevila] (73.8% at 32 captions) and Vamos [wang2023vamos] (72.5%), while having no video level training.
+Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions.
+In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.
 Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks.
@@
 in our cases image domain VLMs and language domain LLMs).
-Table 3: Results on IntentQA [li2023intentqa_cavir]: We compare LVNet against prior methods on IntentQA test set, highlighting standard splits: why?, how?
-and before/after.
-Our method shows state-of-the-art performance in zero-shot, with a significant margin.
-Model Captions Why? (%) How? (%) Before/After (%) All (%)
-finetuned
-HQGA [xiao2022hqga] 16 48.2 54.3 41.7 47.7
-VGT [xiao2022vgt] 32 51.4 56.0 47.6 51.3
-BlindGPT [ouyang2022blindgpt] - 52.2 61.3 43.4 51.6
-CaVIR [li2023intentqa_cavir] - 58.4 65.5 50.5 57.6
-Vamos [wang2023vamos] - 69.5 70.2 65.0 68.5
-zero-shot
-SeViLA [yu2024sevila] 32 - - - 60.9
-LLoVi [zhang2023llovi] 90 68.4 67.4 51.1 64.0
-LangRepo [Kahatapitiya2024langrepo] 90 62.8 62.4 47.8 59.1
-LVNet (ours) 12 75.2 71.6 60.8 71.1
-We finally evaluate on IntentQA [li2023intentqa_cavir] dataset and report these results in Table 3.
-Our approach achieves state-of-the-art results on this benchmark as well outperforming all prior work, both zero-shot and finetuned.
-In fact, LVNet achieves a significant boost (+6.5%) over the closest zero-shot baseline LLoVi [zhang2023llovi] (64.0% at 90 captions), showing a 70.5% accuracy with just 12 frame captions.
-It also extends the state-of-the-art finetuned performance by +2.0%, compared to Vamos [wang2023vamos] (68.5%).
-Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [Mangalam2023EgoSchemaAD].
-We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 280 frames.
-Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM).
-Such coarse keyframes are then ordered according to confidence values, and grouped into 4 groups of 8 frames each.
-These groups go through visal templating (i.e.
-simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
+Qualitative Analysis of the Hierarhical Keyframe Selector:
+We compare the open-ended responses of LVNet and the uniform sampling method in Figure 4 to understand the effectiveness of the hierarchical keyframe selector in LVNet.
+The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively.
+LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames.
+When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, and measuring tape), which is the correct answer.
+This leads LVNet to choose the correct option.
+In contrast, the uniform sampling method selects frames at 0, 16, and 32 seconds and generates captions based on those frames.
+Similarly, when we feed the concatenated captions to the LLM to answer the same question, the output identifies only one activity—welding tools—resulting in the selection of the incorrect option.
+This example highlights the importance of keyframe selection and demonstrates the effectiveness of hierarchical keyframe selection in LVNet.
+Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively.
+LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer.
+The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
 4.3 Ablations
-Table 4: Ablations on EgoSchema [Mangalam2023EgoSchemaAD]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
-Keyframe sampler Seq. captions #captions Acc. (%)
-Uniform Sampling (LLoVi) 180 57.6
-HKS (w/o SVL) 12 63.0
-HKS 12 63.4
+Table 2: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
 Visual Templating order Acc. (%)
-Temporal order 62.9
-Confidence order 63.4
+Temporal 65.2
+Hybrid (Confidence + Temporal) 68.2
+TSC CKD FKD Acc.
+68.2
+65.8
+64.5
+62.6
+Frames 6.4 8 8.4 11 12 16
+VideoAgent[37] 58.4 - 60.2 57.4 - -
+LVNet (ours) - 64.4 - - 68.2 67.8
+In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS.
+In all ablations, we use a subset of EgoSchema [24], composed of 500 videos.
+Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
+Visual Templating Order:
+In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp is more effective than relying solely on temporal order.
+In this hybrid approach, high-confidence frames are selected from shorter segments by sampling three keyframes per set of eight, grouped by confidence rather than timespan.
+Conversely, low-confidence keyframes, which are crucial but more visually challenging for keyword matching, are sampled from broader video sections, with three frames per set of eight grouped by timestamp.
+This hybrid approach outperforms purely temporal ordering by 3%.
+Number of Frame Captions:
+We conducted an ablation study on the number of frame captions used in our setup, comparing it to VideoAgent [37], which operates with a similarly small number of captions.
+Our findings show that LVNet achives its highest accuracy 68.2% with 12 captions, while even its lowest accuracy 64.4% with 8 captions still surpasses VideoAgent’s best accuracy 60.2% with 8.4 captions.
+Effect of Hierarchical Keyframe Modules:
+This table demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules.
+Without any of these modules, the model relies on uniform sampling and achieves 62.6%.
+When TSC is added and 12 frames are selected uniformly, the accuracy increases to 64.5%.
+Adding both TSC and CKD raises the accuracy to 65.8%.
+Finally, incorporating all three modules—TSC, CKD, and FKD—into the model, which is LVNet, results in an accuracy of 68.2%.
+This demonstrates the importance of including all modules in LVNet for optimal performance.
+5 Conclusion
+We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across 3 benchmarks datasets.
+Our Hierarchical Keyframe Selector demonstrates the effectiveness of keyframe selection in understanding a very long-form video QA.
+Additionally, we highlight the zero-shot capability for long-form video comprehension of our LVNet framework, which requires no video-level training.
+Our experiments showcase its significant advantage over previous methods.
+Appendix
+Appendix A.1 Additional Ablations
+In this section, we present additional experiments conducted to inform the LVNet’s design. We have tested different LLMs and experimented with various scales of the visual feature map.
+Table A.3: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
 LLM Acc. (%)
-GPT-3.5 59.2
-GPT-4 61.1
-GPT-4o 63.4
-Frames 6.4 8.4 9 11 12 16
-VideoAgent 58.4 60.2 - 57.4 - -
-LVNet (ours) - - 64.6 - 66.0 65.8
-In this section, we present ablations to key design decisions such as #frames, #words, prompting, and hierarchical processing.
-The results for these ablations are presented in Table 4.
-In all ablations, we consider a subset of 150 videos from EgoSchema [Mangalam2023EgoSchemaAD] for evaluation.
-Keyframe sampler:
-We evaluate our Hierarchical Keyframe Sampling strategy (HKS) against the established uniform sampling baseline from [zhang2023llovi]. Our results clearly indicate the importance of our HKS module.
-Sequential captioning:
-In our SVL module, we consider a sequential captioner, that considers previous captions when captioning the current frame.
-It outperforms the common parallel frame-wise captioning by $+0.4\%$ showing that the SVL can generate better captions, possibly reducing redundant information among captions and enabling a form of temporal modeling.
-Visual Templating order:
-In Visual Templating, grouping the frames based on confidence is better than preserving temporal order.
-It can allow sampling of challenging keyword frames, instead of only selecting high confidence keyframes.
+GPT-3.5 61.0
+GPT-4 65.4
+GPT-4o 68.2
+Patch Size Acc. (%)
+1x1 63.6
+7x7 66.2
+14x14 68.2
 Choice of LLM:
-Here, we see that more-recent GPT-4o outperforms both GPT-4 and GPT-3.5. Since GPT-4o is also cheaper compared to GPT-4, we consider it as our default choice.
-Number of frame captions:
-We ablate different number of frame captions used in our setup compared to VideoAgent [Wang2024VideoAgentLV]. Our method shows more-stable performance.
-5 Conclusion
-We proposed an approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art results across 3 benchmarks datasets.
-Our propsoed Hierarchical Keyframe Selector and Sequential Visual LLM modules demonstrate the usefulness of keyframe selection and sequence-aware frame captioning.
-We also highlight the zero-shot capability for long-form video understanding of our LVNet framework requiring zero video level training, and our experiments demonstrate its clear advantage over prior work.
+Our results show that GPT-4o outperforms both GPT-4 and GPT-3.5 by 2.8% and 7.2%, respectively.
+Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
+Effect of Patch Size on Keyword Matching in CKD:
+Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network [29] to match keywords.
+Our findings show that higher resolutions lead to better accuracy.
+In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
+Appendix A.2 Algorithms in Detail
+Our algorithms are presented in full detail in Algorithm 1, Algorithm 2, and Algorithm 3.
+TSC in Algorithm 1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets.
+Within each of the $n$ sets, we uniformly sample $\leq\tau$ frames, obtaining a total of $T_{a}\leq\tau\times n$ frames.
+For example, LVNet sets $\psi=5,\lambda=12,\tau=18$, resulting in approximately $n\sim 25$ and $T_{a}\sim 390$ on the EgoSchema dataset.
+CKD in Algorithm 2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16.
+LVNet employs $L=32,len(K)\leq 25$ on the EgoSchema dataset.
+FKD in Algorithm 3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally.
+It groups frames sequentially into visual templates, each consisting of $N$ frames.
+From each template, the $M$ frames and keywords most relevant among the $N$ pairs are selected using GPT-4o.
+We set $L=32,K=16,N=8,M=3$.
+1:Require: ResNet-18 [11] pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
+2:for all $img^{i}$ in $\textbf{List}_{frame}$ do
+3:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
+4:   $\textbf{List}_{feat}$.insert($\mathbf{F}^{i}$)
+5:end for
+6:for all $\mathbf{F}^{i}$ in $\textbf{List}_{feat}$ do
+7:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
+8:   $\textbf{M}_{dist}$.insert($\textbf{List}_{dist}$)
+9:end for
+10:while length of $\textbf{List}_{index}>\psi$ do
+11:   $\textbf{List}_{sample}\leftarrow\emptyset$
+12:   $\textbf{List}_{\delta}\leftarrow\emptyset$
+13:   $i\leftarrow\textbf{List}_{index}$.pop(00)
+14:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
+15:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
+16:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
+17:   for all prob in $\textbf{p}^{i}$ do
+18:      if prob < $\beta$ then
+19:         $\textbf{List}_{selected}$.insert($\textbf{index}(\texttt{prob},\textbf{p}^{i})$)
+20:      end if
+21:   end for
+22:   for all $\gamma$ in $\textbf{List}_{selected}$ do
+23:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
+24:      $\textbf{List}_{\delta}$.insert($\delta$)
+25:      $\textbf{List}_{index}$.pop($\gamma$)
+26:   end for
+27:   $\textbf{List}_{\delta}$.insert($i$)
+28:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
+29:   $\textbf{sort}(\textbf{List}_{sample})$
+30:   for all $frame^{j}$ in $\textbf{List}_{frame}$ do
+31:      if $j$ in $\textbf{List}_{sample}$ then
+32:         Outputs.insert($frame^{j}$)
+33:      end if
+34:   end for
+35:end while
+Algorithm 1 Temporal Scene Clustering
+1:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
+2:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
+3:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
+4:Initialize $\mathbf{P}_{best}$ as an empty list
+5:Initialize $\mathbf{I}_{selected}$ as an empty set
+6:while length of $\mathbf{I}_{selected}<L$ do
+7:   for $k\in\mathbf{K}$ do
+8:      for $i\in\mathbf{I}$ do
+9:         $i_{index}\leftarrow\mathbf{idx}_{sorted}[k][i]$
+10:         if $i_{index}$ not in $\mathbf{I}_{selected}$ then
+11:            $\mathbf{P}_{best}$.insert($k,i_{index}$)
+12:            $\mathbf{I}_{selected}$.insert($i_{index}$)
+13:            break
+14:         end if
+15:      end for
+16:      if length of $\mathbf{I}_{selected}\geq L$ then
+17:         break
+18:      end if
+19:   end for
+20:end while
+21:return $\mathbf{P}_{best}$
+Algorithm 2 Keyword-Image Matching Process in CKD
+[…]
+2:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
+[…]
+4:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
+[…]
+6:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
+[…]
+8:for each $\mathbf{set}\in\mathbf{sets}$ do
+9:   $\mathbf{I}_{selected}\leftarrow\texttt{select\_top\_M}(\mathbf{set},M)$
+10:end for
+11:return $\mathbf{I}_{selected}$
+Algorithm 3 Fine Keyframe Detection Process (FKD)
+Appendix A.3 Prompting: Fine Keyframe Detector
+We prompt the VLM to select frames that are most compatible with the list of given keywords.
+Each template image contains 8 images, and their order is described in language (e.g.
+top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure A.5.
+Figure A.5: Prompt for Fine Keyframe Detection: The figure illustrates the input image, the prompt provided to the VLM, and the output.
+The input image represents a visual template composed of eight frames, and the prompt requests the three best frames along with their corresponding keywords.
+The output displays the top three selected frames and their associated keywords.

### v4

@@
 Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector.
-Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, IntentQA, NExT-QA.
+Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, IntentQA.
 The code can be found at https://github.com/jongwoopark7978/LVNet
@@
   • Long-form Video Understanding framework requiring zero video-level training.
-Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], IntentQA[22], and NExT-QA[41] as described in 4.2.
+Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], NExT-QA[41], and IntentQA[22] as described in 4.2.
 This demonstrates the strong performance and general applicability of our approach.
@@
 Datasets are used purely for evaluation.
-We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], IntentQA [22], and NExT-QA [41].
+We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], NExT-QA [41], and IntentQA [22].
 The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions.
 For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip.
-The second dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions.
+The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer.
+These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos.
+The third dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions.
 It consists of 16k multiple-choice questions which are classified as Why?, How?
 or Before/After (B./A.).
-The third dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer.
-These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos.
 Model Choices: For the HKS module, we use the ResNet-18 [12] for the TSC, CLIP-B/16 [29] for the CKD and GPT-4o for the FKD.
@@
 GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework.
-Table 1: Combined Results on IntentQA [22], EgoSchema [24], and NExT-QA [41].
+Table 1: Combined Results on EgoSchema [24], NExT-QA [41], and IntentQA [22].
 We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits.
@@
 Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
-Model IntentQA EgoSchema NExT-QA
-Cap. Why? (%) How? (%) B./A. (%) All (%) Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%)
-Vamos [36] - - - - - - 48.3 - - - - -
-IG-VLM [17] - - - - 65.3 - 59.8 - 69.8 63.6 74.7 68.6
-VideoLLaMA 2 [5] - - - - - - 53.3 - - - - -
-InternVideo2 [38] - - - - - - 60.2 - - - - -
-Tarsier [35] - - - - - - 61.7 - - - - 79.2
-VIOLET [10] - - - - - 5 19.9 - - - - -
-mPLUG-Owl [46] - - - - - 5 31.1 - - - - -
-VideoAgent [37] - - - - - 8.4 54.1 8.2 72.7 64.5 81.1 71.3
-MVU [30] - - - - - 16 37.6 16 55.4 48.1 64.1 55.2
-MoReVQA [25] - - - - - 30 51.7 16 70.2 64.6 - 69.2
-TraveLER [37] - - - - - - - $\sim$25 70.0 60.5 78.2 68.2
-VFC [26] - - - - - - - 32 45.4 51.6 64.1 51.5
-SeViLA [48] 32 - - - 60.9 32 22.7 32 61.3 61.5 75.6 63.6
-ProViQ [6] - - - - - 60 57.1 60 - - - 64.6
-VideoTree [40] $\leq$56 - - - 66.9 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5
-FrozenBiLM [45] - - - - - 90 26.9 - - - - -
-LifelongMemory [39] - - - - - 90 62.1 - - - - -
-LangRepo [16] 90 62.8 62.4 47.8 59.1 180 41.2 90 64.4 51.4 69.1 60.9
-LLoVi [52] 90 68.4 67.4 51.1 64.0 180 50.3 90 69.5 61.0 75.6 67.7
-LVNet (ours) 12 75.0 74.4 62.1 71.7 12 61.1 12 75.0 65.5 81.5 72.9
+Model EgoSchema NExT-QA IntentQA
+Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%) Cap. Why? (%) How? (%) B./A. (%) All (%)
+Vamos [36] - 48.3 - - - - - - - - - -
+IG-VLM [17] - 59.8 - 69.8 63.6 74.7 68.6 - - - - 65.3
+VideoLLaMA 2 [5] - 53.3 - - - - - - - - - -
+InternVideo2 [38] - 60.2 - - - - - - - - - -
+Tarsier [35] - 61.7 - - - - 79.2 - - - - -
+VIOLET [10] 5 19.9 - - - - - - - - - -
+mPLUG-Owl [46] 5 31.1 - - - - - - - - - -
+VideoAgent [37] 8.4 54.1 8.2 72.7 64.5 81.1 71.3 - - - - -
+MVU [30] 16 37.6 16 55.4 48.1 64.1 55.2 - - - - -
+MoReVQA [25] 30 51.7 16 70.2 64.6 - 69.2 - - - - -
+TraveLER [37] - - $\sim$25 70.0 60.5 78.2 68.2 - - - - -
+VFC [26] - - 32 45.4 51.6 64.1 51.5 - - - - -
+SeViLA [48] 32 22.7 32 61.3 61.5 75.6 63.6 32 - - - 60.9
+ProViQ [6] 60 57.1 60 - - - 64.6 - - - - -
+VideoTree [40] 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5 $\leq$56 - - - 66.9
+FrozenBiLM [45] 90 26.9 - - - - - - - - - -
+LifelongMemory [39] 90 62.1 - - - - - - - - - -
+LangRepo [16] 180 41.2 90 64.4 51.4 69.1 60.9 90 62.8 62.4 47.8 59.1
+LLoVi [52] 180 50.3 90 69.5 61.0 75.6 67.7 90 68.4 67.4 51.1 64.0
+LVNet (ours) 12 61.1 12 75.0 65.5 81.5 72.9 12 75.0 74.4 62.1 71.7
 4.2 Evaluation
 Quantitative Results:
-We evaluate LVNet on the IntentQA, EgoSchema, and NExT-QA dataset and present our results in Table 1.
+We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 1.
 Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining.
@@
 The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
-We next evaluate on the IntentQA[22] dataset.
-Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames.
-In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
-We finally evaluate on the NExT-QA[41] dataset.
+We next evaluate on the NExT-QA[41] dataset.
 This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs.
@@
 In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.
+We finally evaluate on the IntentQA[22] dataset.
+Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames.
+In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
 Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks.
@@
 Our experiments showcase its significant advantage over previous methods.
+Acknowledgements:
+This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government [24ZB1200, Research of Human-centered autonomous intelligence system original technology].
+This work was also supported by the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.
+RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 100%)
 Appendix

### v5

@@
 Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames.
-Recent literature explore the use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language.
+Recent literature explore use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language.
 Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant.
 Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector.
-Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, IntentQA.
+Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, and IntentQA, while also demonstrating a strong performance on videos up to an hour long in VideoMME.
 The code can be found at https://github.com/jongwoopark7978/LVNet
+Too Many Frames, Not All Useful:
+Efficient Strategies for Long-Form Video QA
+Jongwoo Park 1  Kanchana Ranasinghe 1  Kumara Kahatapitiya 1 Wonjeong Ryu2  Donghyun Kim 3  Michael S. Ryoo 1 1Stony Brook University   2KAIST AI   3Korea University jongwopark@cs.stonybrook.edu
 1 Introduction
-Video understanding is a long-standing vision problem [1] with numerous real-world applications.
-It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos [3, 15, 33, 13, 32].
-Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos [34, 51, 44].
-The recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures [23, 21, 7], has resulted in significant improvements across visual question answering (VQA) tasks.
-These models demonstrate exceptional performance within the image domain, and their video variants [47, 28, 27] perform similarly on shorter videos, yet demonstrating limited performance on long-form video benchmarks [24, 16, 30, 31].
+Video understanding is a long-standing vision problem (Aggarwal and Ryoo, 2011) with numerous real-world applications.
+It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos (Allen and Ferguson, 1994; Ivanov and Bobick, 2000; Shi et al., 2004; Hongeng et al., 2004; Ryoo and Aggarwal, 2006).
+Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos (Tapaswi et al., 2016; Zeng et al., 2017; Xu et al., 2017).
+Figure 1: LVNet achieves state-of-the-art performance on EgoSchema subset while utilizing only a fraction of captioned frames.
+In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions (VLM calls), outperforming VideoAgent and VideoTree, models using a similar-scale captions, by +8% and +5.7% (more details in Section 4.3).
+Recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures (Liu et al., 2023; Li et al., 2023b; Dai et al., 2023), has resulted in significant improvements across visual question answering (VQA) tasks.
+These models demonstrate exceptional performance within the image domain, and their video variants (Yu et al., 2023; Papalampidi et al., 2023; Maaz et al., 2023) perform similarly on shorter videos, yet demonstrate limited performance on long-form video benchmarks (Mangalam et al., 2023; Kahatapitiya et al., 2024; Rawal et al., 2024).
 This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning.
-An alternate line of works [52, 36, 16, 37] adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain.
+An alternate line of works (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b) adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain.
 However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames.
-Such a design choice leading to high compute expense, is questioned in [4, 30, 37], and is the key motivation for our exploration of key frame selection, i.e.
+Such a design choice leading to high compute expense, is questioned in (Buch et al., 2022; Ranasinghe et al., 2024; Wang et al., 2024b), and is the key motivation for our exploration of key frame selection, i.e.
 identifying a minimal set of frames most useful for correctly answering a given video-question pair.
-Figure 1: (left) Our proposed LVNet contains a novel Hierarchical Keyframe Selector module to select keyframes and followed by VLM and LLM for caption and answer generation.
-Hierarchical Keyframe Selector initially begins by processing dense frames and keywords with the lighter module and progressively exploits heavier and more performance-oriented modules on a small set of frames, resulting from the reduction of keyframe candidates, to ensure efficient computation.
-(right) LVNet achieves state-of-the-art performance on the EgoSchema subset while utilizing only a fraction of captioned frames.
-In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions, outperforming VideoAgent, a model using a similar-scale captions, by 8%.
-A more detailed analysis of accuracy vs.
-the number of captions is provided in 4.3.
-Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Figure 1.
-Aligned with prior work [52, 39, 37], the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair.
+Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Fig.
+2.
+Aligned with prior work (Zhang et al., 2023; Wang et al., 2024d, b), the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair.
+As shown in Fig.
+1, LVNet achieves strong performance using a small set of keyframes from the HKS.
 The scope of this work focuses on optimizing the prior two stages.
 We summarize our key contributions as follows:
-  • Novel Hierarchical Keyframe Selector module which are composed of three submodules: Temporal Scene Clustering (TSC), Coarse Keyframe Detector (CKD), and Fine Keyframe Detector (FKD).
-    • Temporal Scene Clustering performs non-uniform frame sampling by clustering visually similar frames, minimizing redundancy in long videos and capturing key scenes.
-The lightweight module is used to efficiently filter dense frames.
-    • Coarse Keyframe Detector generates keywords, representing atomic activities, using the given query and LLM.
-It assigns confidence scores to frames based on keyword relevance, sub-sampling high-confidence frames for improved interpretability over visual-only selection.
-    • Fine Keyframe Detector refines frame selection by combining multiple frames using visual templating and Visual Language Model (VLM), enabling higher-level reasoning and natural language output for improved accuracy over the CKD’s keyword-based selection.
-  • Long-form Video Understanding framework requiring zero video-level training.
-Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], NExT-QA[41], and IntentQA[22] as described in 4.2.
-This demonstrates the strong performance and general applicability of our approach.
+1. Hierarchical Keyframe Selector (HKS): The proposed HKS consists of three submodules for efficient keyframe selection.
+  1. (a)
+    Temporal Scene Clustering (TSC)
+    • Performs non-uniform frame sampling by clustering visually similar frames.
+    • Reduces redundancy in long videos while capturing key scenes.
+    • A lightweight module for efficient filtering of dense frames.
+  2. (b)
+    Coarse Keyframe Detector (CKD)
+    • Generates keywords representing atomic activities using the given query and an LLM.
+    • Assigns confidence scores to frames based on keyword relevance.
+    • Samples high-confidence frames for improved interpretability over visual-only selection.
+  3. (c)
+    Fine Keyframe Detector (FKD)
+    • Refines frame selection by combining multiple frames using visual templating and a VLM.
+    • Enables higher-level reasoning and natural language-based selection.
+    • Achieves better accuracy than CKD’s keyword-based selection.
+2. Zero-Shot Long-Form Video Understanding: Our framework operates zero-shot without requiring video-level training. This makes it highly efficient for long-form video understanding.
+Figure 2: (top) Overview: LVNet uses a Hierarchical Keyframe Selector (HKS) module to select keyframes, followed by VLM & LLM for caption and answer generation.
+(below) HKS Module processes dense frames with lighter modules and progressively exploits heavier, more performance-oriented modules on smaller subsets of frames to ensure efficient computation.
+Proposed LVNet achieves state-of-the-art results compared to models utilizing similar number of captions on three long-form video question answering benchmarks— EgoSchema, NExT-QA, and IntentQA(Sec.
+4.2).
+This demonstrates strong performance and generality of our approach.
 2 Related Work
-Video Question Answering: Visual question answering (VQA), a relatively recent task, involves generating free-form and open-ended textual content conditioned on an image and natural language text [2].
-Its video variant, Video-VQA [49] replaces images with videos.
-While multiple early datasets focus on querying objects or events based on referential and spatial relations [44, 51, 49], later tasks require explicit temporal understanding of sequential events [18, 19, 50].
-More recent datasets focus on longer videos containing multiple actions and scenes spread over a wide time interval (termed long-form videos) [41, 20] with questions additionally requiring causal reasoning to correctly answer.
-Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning [41] over long temporal windows [24].
-Some works tackling such casual video VQA tasks leverage graph networks to model cross object / event relations [14, 42, 43].
-A more recent line of works integrate large-language models (LLMs) to tackle this task [52, 36, 16, 37, 30, 39, 9] utilizing the strong reasoning skills of the LLMs.
-A common aspect is the use of a vision langauge model (VLM) to convert the frame level visual information into natural language.
-This is in turn input to the LLM which makes a final prediction.
-Unlike these methods, our LVNet incorporates a unique Hierarchical Keyframe Selector that progressively reduces the number of keyframe candidates.
-Lighter modules are applied to dense frames, while heavier, more performance-focused modules are applied to a small set of frames for efficient computation.
+Video Question Answering: Visual question answering (VQA) involves generating open-ended textual content conditioned on an image and natural language query (Agrawal et al., 2015).
+Its video variant, Video-VQA (Yu et al., 2019a) replaces images with videos.
+Multiple early datasets focus on querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019a).
+Later tasks require explicit temporal understanding of sequential events (Lei et al., 2018, 2020; Yu et al., 2019b).
+More recent datasets focus on longer videos containing multiple actions and scenes spread over wide time intervals (termed long-form videos) (Xiao et al., 2021; Li et al., 2022).
+Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023).
+Some works tackling such video VQA tasks leverage graph networks to model cross object / event relations (Hosseini et al., 2022; Xiao et al., 2022a, b).
+A more recent line of works integrate LLMs to tackle this task (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b; Ranasinghe et al., 2024; Wang et al., 2024d; Fan et al., 2024) utilizing the strong reasoning skills of LLMs.
+A common aspect is the use of a vision language model (VLM) to convert frame level visual information into natural language.
+This in turn is input to the LLM which makes a final prediction.
+Unlike these methods, LVNet incorporates a unique Hierarchical Keyframe Selector that progressively reduces the number of keyframe candidates.
+Lighter modules are applied to dense frames, while heavier, more performance-focused modules are applied to a small subset of filtered frames.
 Additionally, LVNet does not require video-level training, unlike earlier supervised approaches.
-Frame Selection in Videos: The task of frame selection in videos has been long explored in video [8, 53] with more recent works focused directly on long-form video question answering [4, 30].
-Most similar to our work is [37] which employs an LLM based strategy for video frame selection.
-However, our proposed Hierarchical Keyframe Selector module differs with unique visual only operations as well as cross-modal operations handling much longer frame sequences extending up to 900 frames.
+Feature Ours VA Tr. VT
+(effective selection)
+Uses non-uniform sampling
+Scene continuity-based selection
+Robust to initial frames
+Fine-grained visual refinement
+(compute efficient)
+Lightweight feature extraction
+Single pass inference
+Table 1: LVNet exhibits unique features compared to VideoAgent (VA) Fan et al. (2024), Traveler (Tr.) Shang et al. (2024) and VideoTree (VT) Wang et al. (2024e). See Appendix A.5 for details.
+Frame Selection in Videos: The task of frame selection in videos has been long explored in video (Davis and Bobick, 1997; Zhao et al., 2017) with more recent works focused directly on long-form video question answering (Buch et al., 2022; Wang et al., 2024e; Fan et al., 2024).
+Most similar to our work is (Wang et al., 2024b) which employs an LLM based strategy for video frame selection.
+However, our LVNet differs with several unique features as summarized in Table 1.
 3 Method
@@
 The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work.
-First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes using ResNet-18, a lightweight module that measures visual correspondence between frames.
-Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords using CLIP-B/16.
+First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords.
 Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM.
@@
 3.1 Background
-Recent approaches utilizing LLMs for long video question answering (LVQA) [52, 36, 16, 30, 37] can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation.
+Recent approaches utilizing LLMs for long video question answering (LVQA) (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Ranasinghe et al., 2024; Wang et al., 2024b) can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation.
 Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g.
@@
 Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
-Figure 2: Detailed architecture of Hierarchical Keyframe Selector: Input frames first go through a Temporal Scene Clustering (TSC) mechanism, which divides the long-video into scenes, enabling per-scene subsampling.
-Next, a Coarse Keyframe Detector (CKD) is applied to select frames best-aligned with keywords relevant to the query.
-Finally, a Fine Keyframe detector further subsamples frames, by refining keyword alignements through a templated visual prompting.
-3.2 Architecture
-Consider a long video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, and width respectively.
-Its paired natural language query is referred as $\mathbf{q}$.
-Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$.
-Our goal is to output the suitable response to this query based on information contained in the video.
-We refer this correct response as $\mathbf{r}$.
-Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$.
-The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$.
-Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$.
-Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
-Our unique contributions here are the Hierarchical Keyframe Selector (HKS) and optimal prompts and description templates for the LLM when generating the final response which is also important as suggested by our experiments.
-The overall architecture is given in Figure 1 (left).
-3.3 Hierarchical Keyframe Selector
-Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [24].
+Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema (Mangalam et al., 2023).
 We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 390 frames.
@@
 simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
+3.2 Architecture
+Consider a video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, width respectively and its paired natural language query $\mathbf{q}$.
+Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$.
+Our goal is to output a response, referred as $\mathbf{r}$, suitable for the given query $\mathbf{q}$ based on information contained in the video $\mathbf{x}$.
+Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$.
+The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$.
+Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$.
+Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
+We illustrate this overall architecture in Figure 2.
+3.3 Hierarchical Keyframe Selector
 We now describe our proposed Hierarchical Keyframe Selector (HKS) module.
-As illustrated in Figure 2, our HKS comprises of three submodules, namely 1) Temporal Scene Clustering (TSC), 2) Coarse Keyframe Detector (CKD), and 3) Fine Keyframe Detector (FKD).
-The inputs to HKS contain $T$ frames with each of the three stages TSC, CKD, and FKD reducing this frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
+As illustrated in Figure 2, our proposed HKS comprises of three sequential submodules, each reducing the frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
 Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling.
-The established approach for preliminary frame selection is uniform sampling.
-In contrast, TSC extracts per-frame visual features using a deep neural network followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets.
+The established approach for preliminary frame selection is uniform sampling (limited to at most 200 frames).
+In contrast, TSC processes 900 to 1800 uniformly sampled frames to extract per-frame visual features using a lightweight deep neural network (ResNet-18) followed by a clustering procedure to identify $n$ non-overlapping frame sets.
 Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$.
 Our iterative clustering procedure is outlined in Algorithm 1.
-It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based grouping operations.
-Our reasoning for TSC is that videos, especially long-form ones, contain extensive information redundancy across its temporal dimension.
-For example, in a long sub-sequence of mostly static frames from a single scene, beyond its first frame the rest may contain little additional information.
-We hypothesize that our TSC module will utilize frame level feature similarity to cluster the frames, such that each cluster corresponds to a common scene or motion.
-A single cluster could contain just one frame or significantly more based on frame feature similarities.
-This leads to a non-uniform sampling of frames across the entire video.
-In contrast to uniform sampling across the entire videos where each timestamp is treated equally, in our formulation the uniform sampling within segments will allow more frames to be sampled from information heavy temporal regions.
-Coarse Keyframe Detector (CKD): Unlike the TSC in prior stage, CKD reasons across both visual and language modalities (using the paired textual query) to further sub-sample the $T_{a}$ frames output from the TSC module.
-The inputs to this CKD stage are the $T_{a}$ frames and the paired query $\mathbf{q}$.
-The module generates a set of keywords $K_{\text{CKD}}$, selects $T_{b}$ frames conditioned on these keywords, and outputs the selected $T_{b}$ frames with a confidence value and keyword assigned to each frame.
-The CKD contains three elements: a keyword generation strategy, a dual-encoder image-text model, and an iterative algorithm for similarity based confidence assignment.
-The keyword generation strategy utilizes the given query, q alongside a hand-crafted templating operation or an LLM to select or generate suitable keywords.
-The dual-encoder image-text model uses a spatially aware contrastive langauge image pre-training (CLIP) network from [29].
+It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based clustering operations.
+A single cluster could contain just one frame or significantly more based on frame feature similarities, leading to a non-uniform sampling of frames across the entire video.
+This allows more frames to be sampled from the information heavy temporal regions of videos.
+Coarse Keyframe Detector (CKD): Unlike TSC in the prior stage, CKD reasons across both visual and language modalities (using the paired textual query, $q$) to further sub-sample $T_{a}$ into $T_{b}$ frames.
+CKD contains three elements: keyword generation strategy, dual-encoder image-text model, and similarity based confidence assignment algorithm.
+Keyword generation utilizes the given query, q, alongside hand-crafted templating operations or an LLM to select or generate suitable keywords.
+The dual-encoder image-text model uses a spatially aware contrastive language image pre-training (CLIP) network from (Ranasinghe et al., 2023).
 For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
-The keyword generation strategy converts the query $\mathbf{q}$ into a set of keywords, $K_{\text{CKD}}$ where the number of keywords $|K_{\text{CKD}}|\in[8,25]$.
-The dual-encoder model converts the keyword set $K_{\text{CKD}}$ and the frame set (containing $T_{a}$ frames) into two sets of embeddings.
-The embeddings of each set are comparable to each other using a cosine similarity distance metric given the training formulation of the CLIP model.
-These two embedding sets are processed by our confidence assignment algorithm in Appendix A.2 which assigns each of the $T_{a}$ frames a confidence score and a keyword.
-Note that the same keyword can be assigned to multiple frames.
-However, given the much higher frame count compared to the keyword count and the cross-frame redundancy, we do not assign multiple keywords to a single frame.
-The key role of this module is language conditioned frame selection.
+See  Appendix A.3 for more details.
 For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query.
@@
 On the visual modality, a single frame will also encode multiple concepts and attributes.
-Our design choice for the spatially aware CLIPpy dual-encoder VLM from [29] is motivated by this nature of individual frames.
-Finally, confidence assignment takes into account these multiple modes of information within each frame and the query and suitably assigns a confidence score to each frame that reflects its relevance to the query.
+Our design choice for the spatially aware CLIPpy dual-encoder VLM from (Ranasinghe et al., 2023) is motivated by this nature of individual frames.
+Finally, confidence assignment takes into account these multiple modes of information within each frame and the query to suitably assign confidence scores to each frame that reflects its query relevance.
 We also highlight how the confidence scores are directly linked to the related keyword (i.e.
@@
 We further illustrate this process in Figure 3 (see Visual Templating column).
-Each of these visual templated images also contain a subset of the keywords that correspond to their 8 images.
+Each of these visual templated images also contain a subset of keywords that correspond to their 8 images.
 These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.4 for details) are input to the VLM.
 The output of the VLM is used to select a subset of each 8 image group.
-These frames are collected as the output of the FKD stage, overall resulting in $T_{c}$ frames.
+These frames are collected as output of the FKD stage, overall resulting in $T_{c}$ frames.
 The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at time).
@@
 Our intuition is that such a mechanism allows one to best utilize the complementary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
-3.4 Comparison with Other Keyframe Selection Methods:
-We aim to highlight the main advantage of the Hierarchical Keyframe Selector over other existing keyframe selection methods.
-Models like VideoAgent and TraveLER provide useful comparisons, as they work with a similar-scale number of captions.
-VideoAgent and TraveLER rely on uniform frame selection in the first iteration without analyzing the entire video.
-They identify important segments based solely on these initial frames and the LLM’s response, which can be problematic if the initial uniformly selected frames are not representative of the entire video or if the LLM misinterprets the captions and prompts.
-In such cases, the LLM might incorrectly identify segments for further analysis.
-If the LLM fails to pinpoint the correct segment initially, the entire process can break down because subsequent frames will be similar to the first set, leading the LLM to continuously select frames within or near the initial segment.
-Additionally, for videos that are as challenging or more difficult than EgoSchema in terms of temporal complexity and activities, these models may require numerous iterations to finalize keyframes selection.
-This results in higher computational and latency costs, as it necessitates numerous runs of resource-intensive VLM and LLM models.
-In contrast, our method analyzes the entire video with high frame rates using a lightweight ResNet-18 [12] and segments the video non-uniformly based on scene continuity.
-We then select several frames in each segment by measuring feature similarity between frame features and keywords using the CLIP-B/16 (0.12B) [29] which is lighted than VideoAgent’s EVA-CLIP-8Bplus (8B).
-By reviewing the entire video and non-uniformly selecting keyframes based on scene continuity and similarity scores, these keyframes accurately represent the question-based important frames distribution in the entire video.
-Furthermore, we use VLM for a fine-grained selection of keyframes, improving keyframe selection when CLIP-B/16 struggles to understand detailed atomic activities in the frames.
-By hierarchically segmenting the video with different modules, the resulting segments and keyframes are more reliable than those from VideoAgent.
-Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.
 4 Experiments
@@
 We then present qualitative results for our method and outline some limitations of our approach.
+Model EgoSchema NExT-QA IntentQA
+Cap. Acc. (%) Cap. Acc. (%) Cap. Acc. (%)
+Vamos (Wang et al., 2023) - 48.3 - - - -
+IG-VLM (Kim et al., 2024) - 59.8 - 68.6 - 65.3
+VideoLLaMA 2 (Cheng et al., 2024) - 53.3 - - - -
+InternVideo2 (Wang et al., 2024c) - 60.2 - - - -
+Tarsier (Wang et al., 2024a) - 61.7 - 79.2 - -
+VIOLET (Fu et al., 2023) 5 19.9 - - - -
+mPLUG-Owl (Ye et al., 2023) 5 31.1 - - - -
+VideoAgent (Wang et al., 2024b) 8.4 54.1 8.2 71.3 - -
+MVU (Ranasinghe et al., 2024) 16 37.6 16 55.2 - -
+MoReVQA (Min et al., 2024) 30 51.7 16 69.2 - -
+VFC (Momeni et al., 2023) - - 32 51.5 - -
+SeViLA (Yu et al., 2024) 32 22.7 32 63.6 32 60.9
+ProViQ (Choudhury et al., 2023) 60 57.1 60 64.6 - -
+VideoTree (Wang et al., 2024e) 62.4 61.1 (56) 73.5 (56) 66.9
+FrozenBiLM (Yang et al., 2022) 90 26.9 - - - -
+LifelongMemory (Wang et al., 2024d) 90 62.1 - - - -
+TraveLER (Shang et al., 2024) (101) 53.3 (65) 68.2 - -
+LangRepo (Kahatapitiya et al., 2024) 180 41.2 90 60.9 90 59.1
+LLoVi (Zhang et al., 2023) 180 50.3 90 67.7 90 64.0
+LVNet (ours) 12 61.1 12 72.9 12 71.7
+Table 2: Long Video Evalation: LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on EgoSchema, NExT-QA, and IntentQA datasets respectively using just 12 frames compared to models using a similar number of captions.
+Models are ordered based on number of captions processed per video.
+Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fair comparison.
+Numbers in parentheses () indicate the maximum number of frames used.
+See Sec.
+A.2 in appendix for detailed results.
+Metric Category VideoAgent VideoTree LVNet
+Avg. Frames $\downarrow$ - 24.6 98.0 24.0
+Acc.(%) $\uparrow$ Knowledge 52.2 60.7 63.0
+Film & TV 42.5 52.5 45.0
+Sports Comp. 42.7 48.6 48.0
+Artistic Perf. 47.5 51.6 53.0
+Life Record 44.7 49.5 45.0
+Multilingual 36.6 40.0 53.0
+[…]
+Average 46.4 53.1 52.4
+Table 3: Evaluating on Very Long Videos.
+Comparison of LVNet (ours) with VideoAgent and VideoTree on the long split of VideoMME.
+LVNet uses the fewest frames while achieving the highest accuracy in three out of six categories and ranking second in overall performance, slightly below the best score.
+In the table, bold indicates the best performance, while underlined represents the second-best performance.
 4.1 Experimental Setup
@@
 Datasets are used purely for evaluation.
-We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], NExT-QA [41], and IntentQA [22].
-The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions.
-For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip.
+We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021), and IntentQA (Li et al., 2023a).
+In addition, to further highlight the strength of our approach on longer videos, we include results on VideoMME’s long splitFu et al.
+(2024).
+These datasets are public available and can be used freely for academic research.
+The first dataset, EgoSchema, consists of 5031 questions and each video lasts three-minute and have multiple choice question.
 The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer.
 These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos.
-The third dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions.
+The third dataset, IntentQA, is based on NExT-QA videos corresponding to temporal and causal reasoning quetions.
 It consists of 16k multiple-choice questions which are classified as Why?, How?
 or Before/After (B./A.).
-Model Choices: For the HKS module, we use the ResNet-18 [12] for the TSC, CLIP-B/16 [29] for the CKD and GPT-4o for the FKD.
+The fourth dataset, VideoMME, consists of very long videos—some up to one hour long, with an average duration of 44 minutes, and provides 900 Q&A.
+Model Choices & Hyperparameters: For the HKS module, we use the ResNet-18 (He et al., 2016a) for the TSC, CLIP-B/16 (Ranasinghe et al., 2023) for the CKD and GPT-4o for the FKD.
 We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T.
 This makes them well-suited for filtering dense frames efficiently.
-In line with previous state-of-the-art work [39, 52, 36], we employ GPT API, especially GPT-4o, for both VLM and LLM.
+In line with previous state-of-the-art work Wang et al.
+(2024d); Zhang et al.
+(2023); Wang et al.
+(2023), we employ GPT API, especially GPT-4o, for both VLM and LLM.
 This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4.
 GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework.
-Table 1: Combined Results on EgoSchema [24], NExT-QA [41], and IntentQA [22].
-We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits.
-The models are ordered based on the number of captions processed before answering the question.
-LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on the three datasets, respectively, using just 12 frames compared to the models using the similar number of captions.
-Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
-Model EgoSchema NExT-QA IntentQA
-Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%) Cap. Why? (%) How? (%) B./A. (%) All (%)
-Vamos [36] - 48.3 - - - - - - - - - -
-IG-VLM [17] - 59.8 - 69.8 63.6 74.7 68.6 - - - - 65.3
-VideoLLaMA 2 [5] - 53.3 - - - - - - - - - -
-InternVideo2 [38] - 60.2 - - - - - - - - - -
-Tarsier [35] - 61.7 - - - - 79.2 - - - - -
-VIOLET [10] 5 19.9 - - - - - - - - - -
-mPLUG-Owl [46] 5 31.1 - - - - - - - - - -
-VideoAgent [37] 8.4 54.1 8.2 72.7 64.5 81.1 71.3 - - - - -
-MVU [30] 16 37.6 16 55.4 48.1 64.1 55.2 - - - - -
-MoReVQA [25] 30 51.7 16 70.2 64.6 - 69.2 - - - - -
-TraveLER [37] - - $\sim$25 70.0 60.5 78.2 68.2 - - - - -
-VFC [26] - - 32 45.4 51.6 64.1 51.5 - - - - -
-SeViLA [48] 32 22.7 32 61.3 61.5 75.6 63.6 32 - - - 60.9
-ProViQ [6] 60 57.1 60 - - - 64.6 - - - - -
-VideoTree [40] 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5 $\leq$56 - - - 66.9
-FrozenBiLM [45] 90 26.9 - - - - - - - - - -
-LifelongMemory [39] 90 62.1 - - - - - - - - - -
-LangRepo [16] 180 41.2 90 64.4 51.4 69.1 60.9 90 62.8 62.4 47.8 59.1
-LLoVi [52] 180 50.3 90 69.5 61.0 75.6 67.7 90 68.4 67.4 51.1 64.0
-LVNet (ours) 12 61.1 12 75.0 65.5 81.5 72.9 12 75.0 74.4 62.1 71.7
+We run TSC and CKD on a single NVIDIA RTX A5000, which takes approximately two hours to process 500 questions.
+We use the default hyperparameters for each vision/language module, as we only perform inference, and set the LLM temperature to 0 to ensure reproducibility.
+Also, We use single run for our experiments.
 4.2 Evaluation
+Model Avg. Frame Captions (Rounded) $\downarrow$
+6 7 8 10 11 12 16 32 62 70
+VideoAgent 58.4 - 63.2 / 60.2 - 60.8 / 57.4 - - - - -
+VideoTree - 61.0 - 61.6 62.2 - - 64.4 66.2 67.0
+LVNet (Ours) - - 64.4 - - 68.2 67.8 - - -
+Templating Order Acc.
+Temporal 65.2
+Confidence 67.6
+Hybrid (both) 68.2
+TSC CKD FKD Acc.
+62.6
+64.5
+65.8
+68.2
+Table 4: Ablation study on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
 Quantitative Results:
@@
 For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions.
-This result outperforms VideoAgent, the next best model using 8.4 captions, by 7%.
-To ensure a fair comparison, we ablate the number of captions for LVNet in LABEL:tab:ablation:numberofframecaptions, observing that LVNet with 8 captions still outperforms VideoAgent with 8.4 captions by 4.2% in the EgoSchema subset.
-The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
-We next evaluate on the NExT-QA[41] dataset.
+This result outperforms VideoAgent, the next best model using 8.4 captions, by +7%, is on par with VideoTree while using only 1/5 of the captions, and outperforms TraveLER by +7.8% while utilizing only 12% of the captions.
+We next evaluate on the NExT-QA dataset.
 This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs.
 Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions.
-In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.
-We finally evaluate on the IntentQA[22] dataset.
-Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames.
-In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
+In fact, our LVNet outperforms VideoAgent by +1.6%.
+In the IntentQA dataset.
+LVNet outperforms all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames.
+In fact, LVNet shows a substantial improvement of +4.8% over the next best model, VideoTree, while using only 13% of the captions (12 vs.
+90).
+Lastly, Table 4 presents the performance of LVNet on VideoMME’s long split, which consists of videos up to one hour long and compare it to other models using keyframe selection methods.
+Our method (LVNet) demonstrates strong performance while utilizing only 24 frames, significantly fewer than VideoTree’s 98 frames.
+LVNet outperforms VideoAgent by +6.0% overall and achieves the highest accuracy in three out of six categories: Knowledge, Artistic Performance, and Multilingual.
+While VideoTree maintains a slight overall lead, LVNet’s ability to achieve comparable accuracy while processing only one-quarter of the frames highlights its efficiency in handling very long videos.
+To ensure a fair comparison, all models utilize GPT-4o.
 Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks.
@@
 in our cases image domain VLMs and language domain LLMs).
-Qualitative Analysis of the Hierarhical Keyframe Selector:
+Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively.
+LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer.
+The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
+Qualitative Analysis of Hierarhical Keyframe Selector:
 We compare the open-ended responses of LVNet and the uniform sampling method in Figure 4 to understand the effectiveness of the hierarchical keyframe selector in LVNet.
@@
 This example highlights the importance of keyframe selection and demonstrates the effectiveness of hierarchical keyframe selection in LVNet.
-Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively.
-LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer.
-The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
 4.3 Ablations
-Table 2: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
-Visual Templating order Acc. (%)
-Temporal 65.2
-Hybrid (Confidence + Temporal) 68.2
-TSC CKD FKD Acc.
-68.2
-65.8
-64.5
-62.6
-Frames 6.4 8 8.4 11 12 16
-VideoAgent[37] 58.4 - 60.2 57.4 - -
-LVNet (ours) - 64.4 - - 68.2 67.8
 In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS.
-In all ablations, we use a subset of EgoSchema [24], composed of 500 videos.
+In all ablations, we use a subset of EgoSchema (Mangalam et al., 2023), composed of 500 videos.
 Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
 Visual Templating Order:
-In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp is more effective than relying solely on temporal order.
-In this hybrid approach, high-confidence frames are selected from shorter segments by sampling three keyframes per set of eight, grouped by confidence rather than timespan.
-Conversely, low-confidence keyframes, which are crucial but more visually challenging for keyword matching, are sampled from broader video sections, with three frames per set of eight grouped by timestamp.
-This hybrid approach outperforms purely temporal ordering by 3%.
+In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp proves more effective than using confidence scores or temporal order alone, as shown in Table LABEL:tab:ablation:visualtemplateorder.
+In this hybrid approach, high-confidence frames capture short but important segments of the video, while low-confidence keyframes, which are crucial but visually challenging for keyword matching, are temporally ordered to cover broader segments.
+This hybrid approach outperforms solely temporal ordering and solely confidence-based ordering by +3% and +0.6%, respectively.
 Number of Frame Captions:
-We conducted an ablation study on the number of frame captions used in our setup, comparing it to VideoAgent [37], which operates with a similarly small number of captions.
-Our findings show that LVNet achives its highest accuracy 68.2% with 12 captions, while even its lowest accuracy 64.4% with 8 captions still surpasses VideoAgent’s best accuracy 60.2% with 8.4 captions.
+We performed an ablation study on the number of frame captions, comparing our approach to VideoAgent (Wang et al., 2024b) and VideoTree (Wang et al., 2024e) under similar low caption settings.
+As shown in Table LABEL:tab:ablation:numberofframecaptions, LVNet achieves the highest accuracy of 68.2% with 12 captions, outperforming VideoAgent (8.4 frames) and VideoTree (12 frames) by +8% and $\sim$+5.7%, respectively.
+We compare LVNet with VideoAgent+GPT-4o (8.1 frames) and VideoTree+GPT-4o (69.5 frames, $\times$5.8 more), both using GPT-4o for a fair comparison.
+We take GPT-4o numbers for VideoAgent and VideoTree from Yang et al.
+(2024) (more details in appendix A.2) and LVNet outperforms them by +5% and +1.2%, respectively.
 Effect of Hierarchical Keyframe Modules:
-This table demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules.
+Table LABEL:tab:ablation:effectofcomponents demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules.
 Without any of these modules, the model relies on uniform sampling and achieves 62.6%.
@@
 Our experiments showcase its significant advantage over previous methods.
+Limitations
+Despite the effectiveness of LVNet, as demonstrated by benchmark experiments and comprehensive ablations, our study has certain limitations, which we discuss below.
+  • First, we acknowledge that we are unable to evaluate LVNet and other models with all available VLMs or LLMs due to computational constraints and high costs.
+However, we carefully select GPT-4o, a state-of-the-art LLM, for our main experiments and provide ablation studies comparing various LLMs (e.g.
+GPT-3.5, GPT-4, and GPT-4o) to other models to ensure a fair performance comparison, as presented in LABEL:tab:ablation:numberofframecaptions and LABEL:tab:ablation:choiceofllm.
+  • Our hierarchical keyframe selector consists of three components: TSC, CKD, and FKD.
+While we demonstrated the effectiveness of each component in LABEL:tab:ablation:effectofcomponents, we did not have the time or resources to develop a unified module that could replace all three.
+Although this is beyond the scope of this paper, exploring a more efficient implementation that integrates these three modules into a single model would be an interesting direction for future research.
+  • Like any LLM-based approach, LVNet is sensitive to prompting.
+To ensure the transparency, we provide examples of these prompts in Figure 4 and Figure A.7.
+We also plan to release the code to enable further exploration by other researchers.
+  • Finally, we acknowledge that, as our approach is zero-shot, any inherent limitations or biases in the pretrained models may persist in the outputs of LVNet.
 Acknowledgements:
@@
 Appendix
+Figure A.5: Comparison of Keyframe Selection: Comparison of LVNet and VideoAgent in keyframe selection for video question answering.
+LVNet refines frames through a multi-stage process (TSC, CKD, FKD) to form a non-uniform keyframe distribution, capturing relevant moments tied to the query.
+In contrast, VideoAgent relies on uniform sampling and LLM-based frame selection, which fails to focus on crucial keyframes, leading to incorrect predictions.
+The final keyframe distributions illustrate LVNet’s ability to retrieve meaningful frames directly related to the answer, while VideoAgent selects irrelevant frames.
 Appendix A.1 Additional Ablations
 In this section, we present additional experiments conducted to inform the LVNet’s design. We have tested different LLMs and experimented with various scales of the visual feature map.
-Table A.3: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
 LLM Acc. (%)
@@
 14x14 68.2
+Table A.5: Additional ablations experiments on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
+Default setting is highlighted.
 Choice of LLM:
-Our results show that GPT-4o outperforms both GPT-4 and GPT-3.5 by 2.8% and 7.2%, respectively.
+Table LABEL:tab:ablation:choiceofllm shows that GPT-4o outperforms both GPT-4 and GPT-3.5 by +2.8% and +7.2%, respectively.
 Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
 Effect of Patch Size on Keyword Matching in CKD:
-Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network [29] to match keywords.
+Table LABEL:tab:ablation:patchsizeinckd shows the effect of the scales of the patch sizes in the CKD.
+Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network (Ranasinghe et al., 2023) to match keywords.
 Our findings show that higher resolutions lead to better accuracy.
 In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
+Table A.6: Exact (Unrounded) Frame Caption Counts. These values supplement the rounded numbers in Table LABEL:tab:ablation:numberofframecaptions All models are based on either GPT-4o or GPT-4.
+Model Avg. Frame Captions $\downarrow$
+6.4 7.1 8 8.1 8.4 9.7 10.7 11 11.3 12 16 32 62.4 69.5
+VideoAgent 58.4 - - 63.2 60.2 - 60.8 57.4 - - - - - -
+VideoTree - 61.0 - - - 61.6 - - 62.2 62.5 66.2 64.4 66.2 67.0
+LVNet (Ours) - - 64.4 - - - - - - 68.2 67.8 - - -
+Exact Frame Caption Counts:
+For completeness, Table A.6 lists the original (unrounded) frame caption counts and corresponding accuracies for VideoAgentWang et al.
+(2024b), VideoTreeWang et al.
+(2024e), and LVNet (ours).
+These values supplement the rounded numbers presented in Table LABEL:tab:ablation:numberofframecaptions of the main text.
+Appendix A.2 Extended results on NExT-QA and IntentQA
+We present extended zero-shot evaluation results on NExT-QA in Table A.7, comparing LVNet with prior zero-shot models across different task categories: causal, temporal, and descriptive reasoning.
+Models are ordered based on the number of captions processed per video, highlighting the trade-offs between caption efficiency and performance.
+LVNet achieves state-of-the-art performance with an overall accuracy of 72.9%, outperforming most models while using only 12 captions per video.
+Notably, it attains 75.0% on causal reasoning, which is the highest among all models evaluated.
+For temporal reasoning, LVNet achieves 65.5%, remaining competitive despite using significantly fewer captions than models like VideoTree (56 captions) and LangRepo (90 captions).
+In descriptive reasoning, LVNet reaches 81.5%, matching VideoTree while processing significantly fewer captions.
+Compared to VideoAgent, the closest competing model in terms of caption efficiency (8.4 captions), LVNet demonstrates a substantial performance gain across all categories, with a +2.8% improvement in overall accuracy.
+While models like VideoTree and TraveLER show strong performance, they process significantly more captions (56 and 65, respectively), indicating that LVNet achieves a superior balance between efficiency and accuracy.
+We present extended zero-shot evaluation results on IntentQA in Table A.8, comparing LVNet with prior zero-shot models across different reasoning categories: Why?, How?, and B.A.
+(Before/After).
+Models are ordered based on the number of captions processed per video, highlighting the balance between caption efficiency and performance.
+LVNet achieves an overall accuracy of 71.7%, outperforming all models while using only 12 captions per video.
+It achieves 75.0% on the Why?
+category, 74.4% on the How?
+category, and 62.1% on the B.A.
+category.
+Compared to VideoTree, which processes 56 captions and achieves an overall accuracy of 66.9%, LVNet outperforms it by +4.8% while using significantly fewer captions.
+Similarly, LangRepo and LLoVi, which process 90 captions, achieve overall scores of 59.1% and 64.0%, respectively, further demonstrating LVNet’s caption efficiency.
+To ensure fairness, models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in grey or downplayed in light green in Table A.7 and Table A.8.
+We adopt the reported GPT-4o results for VideoAgent and VideoTree in Section 4 and LABEL:tab:ablation:numberofframecaptions from the VCA Yang et al.
+(2024), but do not compare directly against VCA for two reasons: (1) the VCA paper does not provide code or implementation details (e.g., inference speed), making replication infeasible; and (2) its reported results cover only a subset of Egoschema, preventing a fair comparison to our approach on the full-scale EgoSchema, NExT-QA, and IntentQA.
+Overeall, these clarifications, alongside the results in Table A.7 and Table A.8, underscore LVNet’s effectiveness in achieving high accuracy while maintaining computational efficiency.
+Model Cap. Cau. (%) Tem. (%) Des. (%) All (%)
+IG-VLM (Kim et al., 2024) - 69.8 63.6 74.7 68.6
+Tarsier (Wang et al., 2024a) - - - - 79.2
+VideoAgent (Wang et al., 2024b) 8.2 72.7 64.5 81.1 71.3
+MVU (Ranasinghe et al., 2024) 16 55.4 48.1 64.1 55.2
+MoReVQA (Min et al., 2024) 16 70.2 64.6 - 69.2
+VFC (Momeni et al., 2023) 32 45.4 51.6 64.1 51.5
+SeViLA (Yu et al., 2024) 32 61.3 61.5 75.6 63.6
+VideoTree (Wang et al., 2024e) (56) 75.2 67.0 81.3 73.5
+ProViQ (Choudhury et al., 2023) 60 - - - 64.6
+TraveLER (Shang et al., 2024) (65) 70.0 60.5 78.2 68.2
+LangRepo (Kahatapitiya et al., 2024) 90 64.4 51.4 69.1 60.9
+LLoVi (Zhang et al., 2023) 90 69.5 61.0 75.6 67.7
+LVNet (ours) 12 75.0 65.5 81.5 72.9
+Table A.7: Extended results on NExT-QA (Xiao et al., 2021).
+We compare LVNet against prior zero-shot models across different reasoning categories: causal, temporal, and descriptive.
+LVNet achieves an overall accuracy of 72.9% while using only 12 captions per video, demonstrating strong performance across all reasoning types.
+Notably, it outperforms all models in causal reasoning (75.0%) and matches the best performance in descriptive reasoning (81.5%), despite processing significantly fewer captions than models like VideoTree (56 captions) and TraveLER (65 captions).
+Models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in gray or downplayed in light green to ensure fairness in comparison.
+Numbers in parentheses () indicate the maximum number of frames used.
+Model Cap. Why? (%) How? (%) B./A. (%) All (%)
+IG-VLM (Kim et al., 2024) - - - - 65.3
+SeViLA (Yu et al., 2024) 32 - - - 60.9
+VideoTree (Wang et al., 2024e) (56) - - - 66.9
+LangRepo (Kahatapitiya et al., 2024) 90 62.8 62.4 47.8 59.1
+LLoVi (Zhang et al., 2023) 90 68.4 67.4 51.1 64.0
+LVNet (ours) 12 75.0 74.4 62.1 71.7
+Table A.8: Extended results on IntentQA (Li et al., 2023a).
+We compare LVNet against prior zero-shot models across different reasoning categories: Why?, How?, and B.A.
+(Belief/Action).
+LVNet achieves an overall accuracy of 71.7%, surpassing all models while using only 12 captions per video.
+It reaches 75.0% in the Why?
+category, 74.4% in the How?
+category, and 62.1% in the B.A.
+category.
+Compared to VideoTree, which processes 56 captions and achieves 66.9% accuracy, LVNet outperforms it by +4.8% while using significantly fewer captions.
+Additionally, LVNet demonstrates superior reasoning-based performance compared to LangRepo (90 captions, 59.1%) and LLoVi (90 captions, 64.0%).
+Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
+Numbers in parentheses () indicate the maximum number of frames used.
 Appendix A.3 Algorithms in Detail
@@
 We set $L=32,K=16,N=8,M=3$.
-1:Require: ResNet-18 [11] pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
+1
+21:Require: ResNet-18 He et al.
+(2016b) pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
+3
 2:for all $img^{i}$ in $\textbf{List}_{frame}$ do
-3:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
+43:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
 4:   $\textbf{List}_{feat}$.insert($\mathbf{F}^{i}$)
-5:end for
+55:end for
 6:for all $\mathbf{F}^{i}$ in $\textbf{List}_{feat}$ do
-7:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
+67:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
 8:   $\textbf{M}_{dist}$.insert($\textbf{List}_{dist}$)
-9:end for
+79:end for
 10:while length of $\textbf{List}_{index}>\psi$ do
@@
 12:   $\textbf{List}_{\delta}\leftarrow\emptyset$
-13:   $i\leftarrow\textbf{List}_{index}$.pop(00)
-14:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
-15:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
-16:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
+813:   $i\leftarrow\textbf{List}_{index}$.pop(00)
+914:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
+1015:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
+1116:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
 17:   for all prob in $\textbf{p}^{i}$ do
@@
 20:      end if
-21:   end for
+1221:   end for
 22:   for all $\gamma$ in $\textbf{List}_{selected}$ do
-23:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
+1323:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
 24:      $\textbf{List}_{\delta}$.insert($\delta$)
 25:      $\textbf{List}_{index}$.pop($\gamma$)
-26:   end for
-27:   $\textbf{List}_{\delta}$.insert($i$)
-28:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
-29:   $\textbf{sort}(\textbf{List}_{sample})$
+1426:   end for
+1527:   $\textbf{List}_{\delta}$.insert($i$)
+1628:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
+1729:   $\textbf{sort}(\textbf{List}_{sample})$
 30:   for all $frame^{j}$ in $\textbf{List}_{frame}$ do
@@
 33:      end if
-34:   end for
+1834:   end for
 35:end while
 Algorithm 1 Temporal Scene Clustering
-1:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
-2:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
-3:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
+1
+21:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
+3
+42:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
+53:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
 4:Initialize $\mathbf{P}_{best}$ as an empty list
-5:Initialize $\mathbf{I}_{selected}$ as an empty set
+65:Initialize $\mathbf{I}_{selected}$ as an empty set
 6:while length of $\mathbf{I}_{selected}<L$ do
@@
 19:   end for
-20:end while
-21:return $\mathbf{P}_{best}$
+720:end while
+821:return $\mathbf{P}_{best}$
 Algorithm 2 Keyword-Image Matching Process in CKD
+1
 […]
-2:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
+3
+42:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
 […]
-4:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
+64:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
 […]
-6:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
+86:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
 […]
@@
 9:   $\mathbf{I}_{selected}\leftarrow\texttt{select\_top\_M}(\mathbf{set},M)$
-10:end for
-11:return $\mathbf{I}_{selected}$
+1010:end for
+$1$:return $\mathbf{I}_{selected}$
 Algorithm 3 Fine Keyframe Detection Process (FKD)
@@
 The output displays the top three selected frames and their associated keywords.
+Appendix A.5 Comparison with Other Keyframe Selection Methods
+We aim to highlight the main advantage of the Hierarchical Keyframe Selector over other existing keyframe selection methods.
+Models like VideoAgent, VideoTree, and TraveLER provide useful comparisons, as they utilize keyframe selection mechanism with similar or different scale of frames.
+VideoAgent and TraveLER rely on uniform frame selection in the first iteration without analyzing the entire video even though they perform non-uniform sampling in the next iterations.
+They identify important segments based solely on these initial frames and the LLM’s response, which can be problematic if the initial uniformly selected frames are not representative of the entire video or if the LLM misinterprets the captions and prompts.
+In such cases, the LLM might incorrectly identify segments for further analysis.
+If the LLM fails to pinpoint the correct segment initially, the entire process can break down because subsequent frames will be similar to the first set, leading the LLM to continuously select frames within or near the initial segment.
+Additionally, for videos that are as challenging or more difficult than EgoSchema in terms of temporal complexity and activities, existing keyframe selection models such as VideoAgent, VideoTree, and TraveLER may require numerous iterations by running heavy visual/language models to finalize keyframes selection.
+This results in higher computational and latency costs, as it necessitates numerous runs of resource-intensive VLM and LLM models.
+In contrast, our method analyzes the entire video with high frame rates using a lightweight ResNet-18 (He et al., 2016a) and segments the video non-uniformly based on scene continuity.
+We then select several frames in each segment by measuring feature similarity between frame features and keywords using the CLIP-B/16 (0.12B) (Ranasinghe et al., 2023) which is lighted than VideoAgent’s EVA-CLIP-8Bplus (8B).
+By reviewing the entire video and non-uniformly selecting keyframes based on scene continuity and similarity scores, these keyframes accurately represent the question-based important frames distribution in the entire video.
+Furthermore, we use VLM for a fine-grained selection of keyframes, improving keyframe selection when CLIP-B/16 struggles to understand detailed atomic activities in the frames.
+By hierarchically segmenting the video with different modules, the resulting segments and keyframes are more reliable than those from VideoAgent.
+Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.
+Figure A.5 visualizes the differences of the keyframe selection mechanism bewtween LVNet and VideoAgent.
+On the left, LVNet begins with uniformly sampled frames and filters them through multiple stages, resulting in a non-uniform distribution of frames over time.
+First, the temporal scene clustering (TSC) selects somes frames that represent temporally distinct activities.
+Next, the coarse keyframe detector (CKD) targets frames most relevant to the question.
+Finally, the fine keyframe detector (FKD) further refines this selection to ensure the keyframes accurately capture the activity in question.
+As a result, LVNet produces 12 frames, with 8 of them (67%) directly depicting "usage of phones," which is the correct answer and leads the model to select the right option.
+On the right, VideoAgent also starts with the uniform frames but relies on a LLM to request additional frames.
+Since the initial frames do not capture enough relevant content, the LLM again selects frames uniformly, adding more irrelevant samples that lack the crucial information about "usage of phones." As a result, VideoAgent ultimately selects the wrong option.

### v6

@@
 Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames.
-Recent literature explore use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language.
+Recent literature leverage large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language.
 Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant.
-Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector.
-Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, and IntentQA, while also demonstrating a strong performance on videos up to an hour long in VideoMME.
+Motivated by this inefficiency, we propose LVNet, a modular and training-free framework featuring a novel Hierarchical Keyframe Selector (HKS) that efficiently selects a minimal set of informative frames tailored to each question.
+LVNet’s modularity allows easy integration with existing approaches for more efficient LVQA.
+We achieve state-of-the-art performance among similarly configured models across four benchmark LVQA datasets: EgoSchema, NExT-QA, IntentQA, VideoMME.
 The code can be found at https://github.com/jongwoopark7978/LVNet
@@
 It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos (Allen and Ferguson, 1994; Ivanov and Bobick, 2000; Shi et al., 2004; Hongeng et al., 2004; Ryoo and Aggarwal, 2006).
-Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos (Tapaswi et al., 2016; Zeng et al., 2017; Xu et al., 2017).
-Figure 1: LVNet achieves state-of-the-art performance on EgoSchema subset while utilizing only a fraction of captioned frames.
-In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions (VLM calls), outperforming VideoAgent and VideoTree, models using a similar-scale captions, by +8% and +5.7% (more details in Section 4.3).
-Recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures (Liu et al., 2023; Li et al., 2023b; Dai et al., 2023), has resulted in significant improvements across visual question answering (VQA) tasks.
-These models demonstrate exceptional performance within the image domain, and their video variants (Yu et al., 2023; Papalampidi et al., 2023; Maaz et al., 2023) perform similarly on shorter videos, yet demonstrate limited performance on long-form video benchmarks (Mangalam et al., 2023; Kahatapitiya et al., 2024; Rawal et al., 2024).
-This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning.
-An alternate line of works (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b) adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain.
-However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames.
-Such a design choice leading to high compute expense, is questioned in (Buch et al., 2022; Ranasinghe et al., 2024; Wang et al., 2024b), and is the key motivation for our exploration of key frame selection, i.e.
+Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering (QnA) that involve generating human-style conversations using large language models (LLMs) (Tapaswi et al., 2016; Zeng et al., 2017; Xu et al., 2017).
+Figure 1: High Accuracy with Low Compute: LVNet achieves state-of-the-art performance on EgoSchema (subset) while processing only a fraction of frame captions (below 12 per video) with the expensive LLM.
+More detailed analysis presented in Section˜$4.$
+Motivated by successful image-language models (Liu et al., 2023; Dai et al., 2023), several works build video LLMs (Yu et al., 2023; Papalampidi et al., 2023; Maaz et al., 2023; Wang et al., 2024b) that directly process video frames in one stage to perform QnA.
+However, for long videos (containing 1000s of frames) these models require processing large visual token sequences with the LLM, making inference computationally expensive or even infeasible (see Table˜5).
+An alternate line of works (Zhang et al., 2023; Wang et al., 2023; Ranasinghe et al., 2024) follow multi-stage pipelines that first extract frame-level information followed by temporal modeling with an LLM.
+While these multi-stage works similarly encounter compute bottlenecks with increased frame processing for longer videos, it is possible to feed the LLM with descriptors of non-uniformly sampled frames Kahatapitiya et al.
+(2024); Wang et al.
+(2024d, f), motivating our exploration into keyframe selection, i.e.
 identifying a minimal set of frames most useful for correctly answering a given video-question pair.
-Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Fig.
-2.
-Aligned with prior work (Zhang et al., 2023; Wang et al., 2024d, b), the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair.
+Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM for caption generation and LLM for answer generation as illustrated in Figure˜2.
+Aligned with prior work (Zhang et al., 2023; Wang et al., 2024f, d), per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair.
 As shown in Fig.
-1, LVNet achieves strong performance using a small set of keyframes from the HKS.
-The scope of this work focuses on optimizing the prior two stages.
+1, LVNet achieves strong performance efficiently, processing only a fraction of frame descriptors (captions) with the LLM.
+Compared to feeding all captions or all frames directly to a powerful model (e.g.
+GPT-4o), our LVNet performs inference at a fraction of the cost (see Table˜6).
 We summarize our key contributions as follows:
-1. Hierarchical Keyframe Selector (HKS): The proposed HKS consists of three submodules for efficient keyframe selection.
   1. (a)
-    Temporal Scene Clustering (TSC)
-    • Performs non-uniform frame sampling by clustering visually similar frames.
-    • Reduces redundancy in long videos while capturing key scenes.
-    • A lightweight module for efficient filtering of dense frames.
+    Efficient Frame Selection: Hierarchical Keyframe Selector (HKS) efficiently extracts keyframes from sequences up to 1800 frames (hour long videos) with a filter rate over 98%.
   2. (b)
-    Coarse Keyframe Detector (CKD)
-    • Generates keywords representing atomic activities using the given query and an LLM.
-    • Assigns confidence scores to frames based on keyword relevance.
-    • Samples high-confidence frames for improved interpretability over visual-only selection.
+    Video Training Free: Our framework requires no video level training and simply uses existing off-the-shelf modules.
   3. (c)
-    Fine Keyframe Detector (FKD)
-    • Refines frame selection by combining multiple frames using visual templating and a VLM.
-    • Enables higher-level reasoning and natural language-based selection.
-    • Achieves better accuracy than CKD’s keyword-based selection.
-2. Zero-Shot Long-Form Video Understanding: Our framework operates zero-shot without requiring video-level training. This makes it highly efficient for long-form video understanding.
+    Versatile Framework: Existing methods can easily be integrated with LVNet to further boost their performance (details in Appendix˜A.7).
 Figure 2: (top) Overview: LVNet uses a Hierarchical Keyframe Selector (HKS) module to select keyframes, followed by VLM & LLM for caption and answer generation.
 (below) HKS Module processes dense frames with lighter modules and progressively exploits heavier, more performance-oriented modules on smaller subsets of frames to ensure efficient computation.
-Proposed LVNet achieves state-of-the-art results compared to models utilizing similar number of captions on three long-form video question answering benchmarks— EgoSchema, NExT-QA, and IntentQA(Sec.
-4.2).
-This demonstrates strong performance and generality of our approach.
+Proposed LVNet achieves state-of-the-art results (at common inference compute) on multiple long-form video question answering benchmarks demonstrating strong performance and generality.
 2 Related Work
@@
 Some works tackling such video VQA tasks leverage graph networks to model cross object / event relations (Hosseini et al., 2022; Xiao et al., 2022a, b).
-A more recent line of works integrate LLMs to tackle this task (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b; Ranasinghe et al., 2024; Wang et al., 2024d; Fan et al., 2024) utilizing the strong reasoning skills of LLMs.
+A more recent line of works integrate LLMs to tackle this task (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024d; Ranasinghe et al., 2024; Wang et al., 2024f; Fan et al., 2024) utilizing the strong reasoning skills of LLMs.
 A common aspect is the use of a vision language model (VLM) to convert frame level visual information into natural language.
@@
 Additionally, LVNet does not require video-level training, unlike earlier supervised approaches.
-Feature Ours VA Tr. VT
+Feature Ours VA Tr. VT VC FV
 (effective selection)
@@
 Single pass inference
-Table 1: LVNet exhibits unique features compared to VideoAgent (VA) Fan et al. (2024), Traveler (Tr.) Shang et al. (2024) and VideoTree (VT) Wang et al. (2024e). See Appendix A.5 for details.
-Frame Selection in Videos: The task of frame selection in videos has been long explored in video (Davis and Bobick, 1997; Zhao et al., 2017) with more recent works focused directly on long-form video question answering (Buch et al., 2022; Wang et al., 2024e; Fan et al., 2024).
-Most similar to our work is (Wang et al., 2024b) which employs an LLM based strategy for video frame selection.
+Video Training Free
+Table 1: LVNet exhibits unique features compared to prior work VideoAgent (VA) Fan et al.
+(2024), Traveler (Tr.) Shang et al.
+(2024), VideoTree (VT) Wang et al.
+(2024g), VideoChat-T (VC) Zeng et al.
+(2024), and Frame-Voyager (FV) Yu et al.
+(2024b) .
+See Appendix˜A.8 for details.
+Frame Selection in Videos: The task of frame selection in videos has been long explored in video (Davis and Bobick, 1997; Zhao et al., 2017) with more recent works focused directly on long-form video question answering (Buch et al., 2022; Wang et al., 2024g; Fan et al., 2024; Zeng et al., 2024; Yu et al., 2024b).
+Most similar to our work is Wang et al.
+(2024d) which employs an LLM based strategy for video frame selection.
 However, our LVNet differs with several unique features as summarized in Table 1.
@@
 Given such a set of useful frames, we also establish optimal strategies for extracting their information using modern large language models (LLMs), taking into account their sequential nature.
-Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure 2.
+Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure˜2.
 The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work.
@@
 3.1 Background
-Recent approaches utilizing LLMs for long video question answering (LVQA) (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Ranasinghe et al., 2024; Wang et al., 2024b) can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation.
+Recent approaches utilizing LLMs for long video question answering (LVQA) (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Ranasinghe et al., 2024; Wang et al., 2024d) can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation.
 Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g.
@@
 Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
-We illustrate this overall architecture in Figure 2.
+We illustrate this overall architecture in Figure˜2.
 3.3 Hierarchical Keyframe Selector
 We now describe our proposed Hierarchical Keyframe Selector (HKS) module.
-As illustrated in Figure 2, our proposed HKS comprises of three sequential submodules, each reducing the frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
+As illustrated in Figure˜2, our proposed HKS comprises of three sequential submodules, each reducing the frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
 Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling.
@@
 Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$.
-Our iterative clustering procedure is outlined in Algorithm 1.
+Our iterative clustering procedure is outlined in Algorithm˜1.
 It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based clustering operations.
@@
 The dual-encoder image-text model uses a spatially aware contrastive language image pre-training (CLIP) network from (Ranasinghe et al., 2023).
-For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
-See  Appendix A.3 for more details.
+For confidence assignment, we construct an algorithm as outlined in Algorithm˜2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
+See  Appendix˜A.5 for more details.
 For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query.
@@
 Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame.
-We further illustrate this process in Figure 3 (see Visual Templating column).
+We further illustrate this process in Figure˜3 (see Visual Templating column).
 Each of these visual templated images also contain a subset of keywords that correspond to their 8 images.
-These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.4 for details) are input to the VLM.
+These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix˜A.6 for details) are input to the VLM.
 The output of the VLM is used to select a subset of each 8 image group.
@@
 A total of 16 low-score frames are temporally reordered in this process.
-The algorithm is described in Algorithm 3 and the prompting technique is explained in Appendix A.4.
+The algorithm is described in Algorithm˜3 and the prompting technique is explained in Appendix˜A.6.
 Our intuition is that such a mechanism allows one to best utilize the complementary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
@@
 In this section, we first discuss our experimental setup followed by quantitative evaluations comparing to existing baselines and ablations of our proposed components.
-We then present qualitative results for our method and outline some limitations of our approach.
-Model EgoSchema NExT-QA IntentQA
-Cap. Acc. (%) Cap. Acc. (%) Cap. Acc. (%)
-Vamos (Wang et al., 2023) - 48.3 - - - -
-IG-VLM (Kim et al., 2024) - 59.8 - 68.6 - 65.3
-VideoLLaMA 2 (Cheng et al., 2024) - 53.3 - - - -
-InternVideo2 (Wang et al., 2024c) - 60.2 - - - -
-Tarsier (Wang et al., 2024a) - 61.7 - 79.2 - -
-VIOLET (Fu et al., 2023) 5 19.9 - - - -
-mPLUG-Owl (Ye et al., 2023) 5 31.1 - - - -
-VideoAgent (Wang et al., 2024b) 8.4 54.1 8.2 71.3 - -
-MVU (Ranasinghe et al., 2024) 16 37.6 16 55.2 - -
-MoReVQA (Min et al., 2024) 30 51.7 16 69.2 - -
-VFC (Momeni et al., 2023) - - 32 51.5 - -
-SeViLA (Yu et al., 2024) 32 22.7 32 63.6 32 60.9
-ProViQ (Choudhury et al., 2023) 60 57.1 60 64.6 - -
-VideoTree (Wang et al., 2024e) 62.4 61.1 (56) 73.5 (56) 66.9
-FrozenBiLM (Yang et al., 2022) 90 26.9 - - - -
-LifelongMemory (Wang et al., 2024d) 90 62.1 - - - -
-TraveLER (Shang et al., 2024) (101) 53.3 (65) 68.2 - -
-LangRepo (Kahatapitiya et al., 2024) 180 41.2 90 60.9 90 59.1
-LLoVi (Zhang et al., 2023) 180 50.3 90 67.7 90 64.0
-LVNet (ours) 12 61.1 12 72.9 12 71.7
-Table 2: Long Video Evalation: LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on EgoSchema, NExT-QA, and IntentQA datasets respectively using just 12 frames compared to models using a similar number of captions.
+We then discuss the results on efficiency and scalability, and finally outline limitations.
+Model EgoSchema NExT-QA IntentQA VT Free
+[…]
+VideoLLaMA 2 (Cheng et al., 2024) - 53.3 - - - - no
+InternVideo2 (Wang et al., 2024e) - 60.2 - - - - no
+Tarsier (Wang et al., 2024a) - 61.7 - 79.2 - - no
+Vamos (Wang et al., 2023) - 48.3 - - - - yes
+IG-VLM (Kim et al., 2024) - 59.8 - 68.6 - 65.3 yes
+VIOLET (Fu et al., 2023) 5 19.9 - - - - yes
+mPLUG-Owl (Ye et al., 2023) 5 31.1 - - - - yes
+VideoAgent (Wang et al., 2024d) 8.4 54.1 8.2 71.3 - - yes
+MVU (Ranasinghe et al., 2024) 16 37.6 16 55.2 - - yes
+MoReVQA (Min et al., 2024) 30 51.7 16 69.2 - - yes
+VFC (Momeni et al., 2023) - - 32 51.5 - - yes
+ProViQ (Choudhury et al., 2023) 50 57.1 50 64.6 - - yes
+VideoTree (Wang et al., 2024g) 62.4 61.1 (56) 73.5 (56) 66.9 yes
+FrozenBiLM (Yang et al., 2022) 90 26.9 - - - - yes
+LifelongMemory (Wang et al., 2024f) 90 62.1 - - - - yes
+TraveLER (Shang et al., 2024) (101) 53.3 (65) 68.2 - - yes
+LangRepo (Kahatapitiya et al., 2024) 180 41.2 90 60.9 90 59.1 yes
+LLoVi (Zhang et al., 2023) 180 50.3 90 67.7 90 64.0 yes
+LVNet (ours) 12 61.1 12 72.9 12 71.7 yes
+Table 2: Long Video Evalation: LVNet achieves state-of-the-art accuracies of 61.1%, 72.9%, and 71.7% on EgoSchema, NExT-QA, and IntentQA datasets respectively using just 12 frames compared to models using a similar number of captions.
 Models are ordered based on number of captions processed per video.
-Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fair comparison.
+Models with video-level training (VT) or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fair comparison.
 Numbers in parentheses () indicate the maximum number of frames used.
@@
 A.2 in appendix for detailed results.
-Metric Category VideoAgent VideoTree LVNet
-Avg. Frames $\downarrow$ - 24.6 98.0 24.0
-Acc.(%) $\uparrow$ Knowledge 52.2 60.7 63.0
-Film & TV 42.5 52.5 45.0
-Sports Comp. 42.7 48.6 48.0
-Artistic Perf. 47.5 51.6 53.0
-Life Record 44.7 49.5 45.0
-Multilingual 36.6 40.0 53.0
 […]
-Average 46.4 53.1 52.4
-Table 3: Evaluating on Very Long Videos.
-Comparison of LVNet (ours) with VideoAgent and VideoTree on the long split of VideoMME.
-LVNet uses the fewest frames while achieving the highest accuracy in three out of six categories and ranking second in overall performance, slightly below the best score.
-In the table, bold indicates the best performance, while underlined represents the second-best performance.
+VideoChat-T 7B/OS no N/A 43.8
+Frame-Voyager 34B/OS no N/A 51.2
+LVNet (DS-V3) 37B/OS yes 24 53.1
+VideoAgent+GPT-4o <1.8T/$P$ yes 24 51.3
+VideoTree+GPT-4o <1.8T/$P$ yes 24 52.3
+LVNet (GPT-4o) <1.8T/$P$ yes 24 53.9
+Table 3: Comparison to Keyframe Selection Methods on VideoMME: We compare LVNet with both single-stage methods that rely on video-level training (VideoChat-T, Frame-Voyager) and two-stage methods (VideoAgent, VideoTree).
+For the two-stage comparison, we evaluate VideoTree and VideoAgent with GPT-4o as both the captioner and LLM under the same 24-frame caption budget for a direct head-to-head comparison.
+Legend: DS-V3=DeepSeek-V3, OS=open-source, PP=proprietary, VT Free=no video-level training, TS Cap.=two-stage caption numbers..
+Notably, LVNet outperforms both single-stage and two-stage methods without requiring video-level training.
 4.1 Experimental Setup
@@
 Datasets are used purely for evaluation.
-We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021), and IntentQA (Li et al., 2023a).
-In addition, to further highlight the strength of our approach on longer videos, we include results on VideoMME’s long splitFu et al.
-(2024).
+We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021), and IntentQA (Li et al., 2023).
+In addition, to further highlight the strength of our approach on longer videos, we include results on VideoMME’s long split (Fu et al., 2024).
 These datasets are public available and can be used freely for academic research.
@@
 The fourth dataset, VideoMME, consists of very long videos—some up to one hour long, with an average duration of 44 minutes, and provides 900 Q&A.
+Collectively, these benchmarks span three orders of magnitude in duration—from short clips ($\sim$44 s on NExT‑QA/IntentQA), to minute‑scale videos (EgoSchema, $\approx$3 min), to hour‑scale footage (VideoMME’s long split, 30–60 min)—allowing us to stress‑test LVNet from seconds to hour‑long videos under a single, training‑free evaluation protocol.
 Model Choices & Hyperparameters: For the HKS module, we use the ResNet-18 (He et al., 2016a) for the TSC, CLIP-B/16 (Ranasinghe et al., 2023) for the CKD and GPT-4o for the FKD.
-We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T.
-This makes them well-suited for filtering dense frames efficiently.
+We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to LLMs that are on a billion parameter scale.
 In line with previous state-of-the-art work Wang et al.
-(2024d); Zhang et al.
+(2024f); Zhang et al.
 (2023); Wang et al.
-(2023), we employ GPT API, especially GPT-4o, for both VLM and LLM.
-This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4.
-GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework.
-We run TSC and CKD on a single NVIDIA RTX A5000, which takes approximately two hours to process 500 questions.
-We use the default hyperparameters for each vision/language module, as we only perform inference, and set the LLM temperature to 0 to ensure reproducibility.
-Also, We use single run for our experiments.
+(2023), we use up to 4 NVIDIA RTX A5000 GPUs and GPT or DeepSeek APIs for running baselines and our setup in all evaluations.
+Also following prior work, we report results over single evaluation runs.
 4.2 Evaluation
-Model Avg. Frame Captions (Rounded) $\downarrow$
-6 7 8 10 11 12 16 32 62 70
-VideoAgent 58.4 - 63.2 / 60.2 - 60.8 / 57.4 - - - - -
-VideoTree - 61.0 - 61.6 62.2 - - 64.4 66.2 67.0
-LVNet (Ours) - - 64.4 - - 68.2 67.8 - - -
-Templating Order Acc.
+Model Frame Captions
+8 12 16
+VideoAgent+GPT-4o 63.4 63.6 64.6
+VideoTree+GPT-4o 62.8 63.6 66.4
+LVNet (GPT-4o) 64.4 68.2 67.8
+Templating Order Acc. $\uparrow$
 Temporal 65.2
@@
 Hybrid (both) 68.2
-TSC CKD FKD Acc.
+TSC CKD FKD Acc. $\uparrow$
 62.6
@@
 Table 4: Ablation study on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
+Method LLM AP Frames $\uparrow$ LLM frames $\downarrow$ Acc. $\uparrow$
+Qwen-VL 7B 128 128 37.8
+Qwen-VL 7B 256 256 OOM
+Qwen-VL 7B 1800 1800 OOM
+LVNet (DS-V3) 37B 1800 24 53.1
+Table 5: Single-Stage Method Comparison: We report Accuracy on VideoMME long split (average video length 41 mins) along with LLM active parameters (LLM AP), total frames processed, and frames input to LLM as tokens / captions (LLM frames).
+LVNet DeepSeek-V3 (DS-V3) variant is used.
+Processing lengthy frame sequences with single-stage models at fixed compute becomes infeasible.
+Inference tested on 4 x 24GB A5000 GPUs.
+Similar findings in Zeng et al.
+(2024).
 Quantitative Results:
@@
 Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency.
-For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions.
+We reiterate how number of captions input to LLM affects inference compute cost of methods the most.
+For EgoSchema(fullset), LVNet achieves 61.1%, the highest among the models utilizing approximately 12 captions.
 This result outperforms VideoAgent, the next best model using 8.4 captions, by +7%, is on par with VideoTree while using only 1/5 of the captions, and outperforms TraveLER by +7.8% while utilizing only 12% of the captions.
-We next evaluate on the NExT-QA dataset.
-This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs.
-Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions.
+We next evaluate on NExT-QA dataset, which has a particular focus on both temporal and casual reasoning based question-answer pairs.
+LVNet achieves state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions.
 In fact, our LVNet outperforms VideoAgent by +1.6%.
-In the IntentQA dataset.
-LVNet outperforms all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames.
-In fact, LVNet shows a substantial improvement of +4.8% over the next best model, VideoTree, while using only 13% of the captions (12 vs.
-90).
-Lastly, Table 4 presents the performance of LVNet on VideoMME’s long split, which consists of videos up to one hour long and compare it to other models using keyframe selection methods.
-Our method (LVNet) demonstrates strong performance while utilizing only 24 frames, significantly fewer than VideoTree’s 98 frames.
-LVNet outperforms VideoAgent by +6.0% overall and achieves the highest accuracy in three out of six categories: Knowledge, Artistic Performance, and Multilingual.
-While VideoTree maintains a slight overall lead, LVNet’s ability to achieve comparable accuracy while processing only one-quarter of the frames highlights its efficiency in handling very long videos.
-To ensure a fair comparison, all models utilize GPT-4o.
-Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks.
-Nevertheless, we highlight how our zero-shot and video level training-free framework is competitive with the best supervised approaches on this dataset.
-This indicates the promise of utilizing pretrained models, especially those equipped with extensive world knowledge and reasoning skills from alternate modality specific learning (i.e.
-in our cases image domain VLMs and language domain LLMs).
-Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively.
-LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer.
-The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
-Qualitative Analysis of Hierarhical Keyframe Selector:
-We compare the open-ended responses of LVNet and the uniform sampling method in Figure 4 to understand the effectiveness of the hierarchical keyframe selector in LVNet.
-The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively.
-LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames.
-When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, and measuring tape), which is the correct answer.
-This leads LVNet to choose the correct option.
-In contrast, the uniform sampling method selects frames at 0, 16, and 32 seconds and generates captions based on those frames.
-Similarly, when we feed the concatenated captions to the LLM to answer the same question, the output identifies only one activity—welding tools—resulting in the selection of the incorrect option.
-This example highlights the importance of keyframe selection and demonstrates the effectiveness of hierarchical keyframe selection in LVNet.
+In IntentQA dataset, LVNet outperforms all prior work, including de-emphasized models with video-caption pretraining and downplayed models utilizing significantly compute (captions input to LLM) than ours.
+In fact, LVNet shows a substantial improvement of +4.8% over the next best VideoTree, while using only 13% of captions (12 vs 90).
+Lastly, Table 3 shows LVNet’s performance on VideoMME’s long split (Fu et al., 2024), which features videos up to an hour long—significantly exceeding the 12-minute average in MLVU (Zhou et al., 2024) or the 17-minute average in the overall VideoMME.
+In combination with an open-source LLM (DeepSeek-V3), LVNet outperforms the single-stage keyframe selection methods VideoChat-T (Zeng et al., 2024) and Frame-Voyager (Yu et al., 2024b) by +9.3% and +1.9%, respectively, without any video-level training.
+Moreover, under a direct head-to-head setting where VideoAgent, VideoTree, and LVNet all use GPT-4o as both the frame captioner and LLM with the same 24-frame budget, LVNet achieves 53.9% accuracy, outperforming VideoAgent(51.3%) and VideoTree(52.3%) by +2.6% and +1.6%, respectively, further confirming its effectiveness on very long videos.
+See Appendix˜A.3 for the detailed results.
+Qualitative Examples:
+Due to space, qualitative open-ended responses comparing LVNet to uniform sampling are provided in the Appendix (see Figure˜A.6).
+Model FV $\uparrow$ FL $\downarrow$ VC $\downarrow$ TC $\downarrow$ Acc. $\uparrow$
+GPT-4o 384 384 \$2.88 $\sim$\$2592 65.3
+LVNet (GPT-4o) 1800 24 \$0.19 \$171 53.9
+Table 6: API Comparison: We perform cost comparison with using GPT-4o directly on frames vs with LVNet.
+We report accuracy on VideoMME long split (Acc), frames processed per video (FV), frames / captions input to LLM (FL), per video cost (VC), and total evaluation cost (TC).
+Our LVNet achieves competitive performance at over 10x less inference cost.
+GPT-4o accuracy result from official benchmark leaderboard.
+4.3 Efficiency & Scalability
+[…]
+Stages FPS Mem. Frames Lat. (s) Lat. (%)
+HKS-1/3 (TSC) 182 19 1800 9.9 16
+HKS-2/3 (CKD) 138 9 427 3.1 5
+HKS-3/3 (FKD) 1 - 8 9.2 15
+Captioning 1 - 24 38.7 62
+LLM Q&A - - - 1.0 2
+Table 7: LVNet Stage-wise Performance: The table reports per-module elapsed time, FPS, and peak GPU memory for a single VideoMME-long video processed on an RTX-6000 Ada.
+Within LVNet, the three stage Hierarchical Keyframe Selectors contributes 36 % of the total runtime.
+It indicates that most compute lies in the captioning and LLM Q&A steps, which are overheads shared by all two-stage pipelines, rather than in the keyframe selection.
+Mem.
+and Lat.
+in the header stands for the GPU memory (GB) and Latency.
+[…]
+[…]
+VideoAgent 87.3 2.1
+VideoTree 64.2 2.8
+LVNet 25.6 7.0
+[…]
+[…]
+Model Frames Memory (GB) Inference time (s) FPS
+Qwen2.5-VL-72B 8 355 11.8 0.7
+Qwen2.5-VL-72B 12 373 16.8 0.7
+Qwen2.5-VL-72B 24 OOM OOM OOM
+LVNet 1800 19 61.9 29.1
+Table 9: Model Efficiency Comparison: End-to-end, LVNet handles 1,800 frames in 61.9 s (29.1 FPS) while using just 19 GB on a single GPU, whereas Qwen2.5-VL-72B requires eight GPUs and 373 GB yet runs $\approx 40\times$ slower at 0.7 FPS.
+Despite architectural and hardware differences, these figures show that LVNet keyframe selector is lightweight and that the overall system delivers high FPS while remaining single-GPU friendly for long-form video.
+Single-stage scaling limits on long videos.
+Table 5 shows that single-stage models (e.g., Qwen-VL (Bai et al., 2023)) run OOM beyond 128 frames, whereas LVNet achieves higher accuracy with only 24 keyframes passed to the LLM, maintaining tractable compute.
+It shows the LVNet’s scalability of processing large number of frames with high accuracy.
+[…]
+Scalability across frame budget and LLM upgrades.
+Table˜10 summarizes two complementary scaling axes.
+Left: on VideoMME‑Long, increasing the frame budget from 12 to 24 yields a +3.8% gain.
+Conversely, on EgoSchema’s 3-min videos, the 12-frame setting is optimal (68.2 %), indicating extra keyframes help LVNet for longer videos.
+It is consistent with our keyframe‑driven design that benefits from additional informative frames when videos are very long.
+Right: LVNet improves monotonically with stronger LLMs (GPT‑3.5 $\rightarrow$ GPT‑4 $\rightarrow$ GPT‑4o) without any retraining, highlighting the modularity and scalability of LVNet with newer LLMs.
+Frames Acc. (%)
+12 50.1
+24 53.9   (+3.8)
+LLM Acc. (%)
+GPT-3.5 61.0
+GPT-4 65.4
+GPT-4o 68.2
+Table 10: Scalability of LVNet. Left: accuracy improves as the frame budget increases for very long videos. Right: newer LLMs yield higher accuracy without retraining, reflecting LVNet’s modularity.
 4.4 Ablations
@@
 In all ablations, we use a subset of EgoSchema (Mangalam et al., 2023), composed of 500 videos.
-Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
+Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix˜A.1
+Number of Frame Captions:
+We conduct an ablation on the number of frame captions, directly comparing our approach with VideoAgent (Wang et al., 2024d) and VideoTree (Wang et al., 2024g) under the same GPT-4o backbone for both frame captioning and LLM reasoning.
+As shown in Table LABEL:tab:ablation:numberofframecaptions, LVNet achieves the highest accuracy of 68.2% with 12 captions, outperforming both VideoAgent and VideoTree (63.6%) by +4.6% absolute.
+LVNet also consistently outperforms the baselines at 8 captions (+1.0% / +1.6%) and 16 captions (+3.2% / +1.4%) compared to VideoAgent / VideoTree..
 Visual Templating Order:
 In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp proves more effective than using confidence scores or temporal order alone, as shown in Table LABEL:tab:ablation:visualtemplateorder.
-In this hybrid approach, high-confidence frames capture short but important segments of the video, while low-confidence keyframes, which are crucial but visually challenging for keyword matching, are temporally ordered to cover broader segments.
-This hybrid approach outperforms solely temporal ordering and solely confidence-based ordering by +3% and +0.6%, respectively.
-Number of Frame Captions:
-We performed an ablation study on the number of frame captions, comparing our approach to VideoAgent (Wang et al., 2024b) and VideoTree (Wang et al., 2024e) under similar low caption settings.
-As shown in Table LABEL:tab:ablation:numberofframecaptions, LVNet achieves the highest accuracy of 68.2% with 12 captions, outperforming VideoAgent (8.4 frames) and VideoTree (12 frames) by +8% and $\sim$+5.7%, respectively.
-We compare LVNet with VideoAgent+GPT-4o (8.1 frames) and VideoTree+GPT-4o (69.5 frames, $\times$5.8 more), both using GPT-4o for a fair comparison.
-We take GPT-4o numbers for VideoAgent and VideoTree from Yang et al.
-(2024) (more details in appendix A.2) and LVNet outperforms them by +5% and +1.2%, respectively.
+In this hybrid approach, high-confidence frames capture short but important segments of videos, while low-confidence keyframes, which are crucial but visually challenging for keyword matching, are temporally ordered to cover broader segments.
+This hybrid approach outperforms solely temporal ordering or confidence-based ordering by +3% and +0.6%, respectively.
 Effect of Hierarchical Keyframe Modules:
@@
 5 Conclusion
-We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across 3 benchmarks datasets.
+We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across four benchmarks.
 Our Hierarchical Keyframe Selector demonstrates the effectiveness of keyframe selection in understanding a very long-form video QA.
 Additionally, we highlight the zero-shot capability for long-form video comprehension of our LVNet framework, which requires no video-level training.
-Our experiments showcase its significant advantage over previous methods.
+Our experiments showcase its significant advantage over existing single-stage and two-stage keyframe selectors.
 Limitations
@@
   • First, we acknowledge that we are unable to evaluate LVNet and other models with all available VLMs or LLMs due to computational constraints and high costs.
-However, we carefully select GPT-4o, a state-of-the-art LLM, for our main experiments and provide ablation studies comparing various LLMs (e.g.
-GPT-3.5, GPT-4, and GPT-4o) to other models to ensure a fair performance comparison, as presented in LABEL:tab:ablation:numberofframecaptions and LABEL:tab:ablation:choiceofllm.
+However, we carefully select GPT-4o and DeepSeek-v3, the LLMs commonly used in video understanding research, for our main experiments and provide ablation studies comparing various LLMs (e.g.
+GPT-3.5, GPT-4, and GPT-4o) to ensure a fair performance comparison, as presented in LABEL:tab:ablation:numberofframecaptions and LABEL:tab:llm_choice_eff.
   • Our hierarchical keyframe selector consists of three components: TSC, CKD, and FKD.
@@
   • Like any LLM-based approach, LVNet is sensitive to prompting.
-To ensure the transparency, we provide examples of these prompts in Figure 4 and Figure A.7.
+To ensure the transparency, we provide examples of these prompts in Figure˜A.6 and Figure˜A.7.
 We also plan to release the code to enable further exploration by other researchers.
@@
 Acknowledgements:
-This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government [24ZB1200, Research of Human-centered autonomous intelligence system original technology].
-This work was also supported by the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.
-RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 100%)
+This research was financially supported by the Ministry of Trade, Industry, and Energy (MOTIE), Korea, under the “Global Industrial Technology Cooperation Center(GITCC) program” supervised by the Korea Institute for Advancement of Technology (KIAT).(Task No.
+P0028420).
+This research was supported by the National Research Council of Science & Technology(NST) grant by the Korea government(MSIT) (No.
+GTL25041-000).
 Appendix
@@
 In this section, we present additional experiments conducted to inform the LVNet’s design. We have tested different LLMs and experimented with various scales of the visual feature map.
-LLM Acc. (%)
-GPT-3.5 61.0
-GPT-4 65.4
-GPT-4o 68.2
+Model Frames Acc. (%)
+GPT-4o 12 62.6
+LVNet 8 64.4
+LVNet 12 68.2
 Patch Size Acc. (%)
@@
 14x14 68.2
-Table A.5: Additional ablations experiments on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
-Default setting is highlighted.
-Choice of LLM:
-Table LABEL:tab:ablation:choiceofllm shows that GPT-4o outperforms both GPT-4 and GPT-3.5 by +2.8% and +7.2%, respectively.
-Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
+Table A.11: Additional comparisons on EgoSchema subset. Left: uniform 12‑frame sampling (GPT‑4o) vs LVNet at 8 and 12 frames. Right: effect of patch size in CKD.
+Uniform Sampling Comparison:
+GPT-4o with uniform 12‑frame sampling on the EgoSchema subset scores 62.6%.
+LVNet outperforms GPT-4o, reaching 64.4% with just 8 frames and 68.2% with 12 frames (Table LABEL:tab:appendix_uniform).
+To keep the evaluation both methodical and cost‑effective, we first ran GPT-4o and LVNet on the EgoSchema subset as validation to find the optimal settings, then transferred those settings to the full EgoSchema test set and the remaining benchmarks.
+This procedure avoids the $\mathdollar 3{,}000$ API expense of running GPT-4o uniformly across all four datasets while still demonstrating LVNet’s stronger generalization.
+For LLM choices, see Sec.
+4.3.
 Effect of Patch Size on Keyword Matching in CKD:
@@
 Our findings show that higher resolutions lead to better accuracy.
-In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
-Table A.6: Exact (Unrounded) Frame Caption Counts. These values supplement the rounded numbers in Table LABEL:tab:ablation:numberofframecaptions All models are based on either GPT-4o or GPT-4.
-Model Avg. Frame Captions $\downarrow$
-6.4 7.1 8 8.1 8.4 9.7 10.7 11 11.3 12 16 32 62.4 69.5
-VideoAgent 58.4 - - 63.2 60.2 - 60.8 57.4 - - - - - -
-VideoTree - 61.0 - - - 61.6 - - 62.2 62.5 66.2 64.4 66.2 67.0
-LVNet (Ours) - - 64.4 - - - - - - 68.2 67.8 - - -
-Exact Frame Caption Counts:
-For completeness, Table A.6 lists the original (unrounded) frame caption counts and corresponding accuracies for VideoAgentWang et al.
-(2024b), VideoTreeWang et al.
-(2024e), and LVNet (ours).
-These values supplement the rounded numbers presented in Table LABEL:tab:ablation:numberofframecaptions of the main text.
+In LVNet, we use a $\times14$ feature map and determine the confidence level of the keyword by selecting the maximum value between the $\times14$ patches and the keyword’s text embedding.
 Appendix A.2 Extended results on NExT-QA and IntentQA
-We present extended zero-shot evaluation results on NExT-QA in Table A.7, comparing LVNet with prior zero-shot models across different task categories: causal, temporal, and descriptive reasoning.
+We present extended zero-shot evaluation results on NExT-QA in Table˜A.12, comparing LVNet with prior zero-shot models across different task categories: causal, temporal, and descriptive reasoning.
 Models are ordered based on the number of captions processed per video, highlighting the trade-offs between caption efficiency and performance.
@@
 While models like VideoTree and TraveLER show strong performance, they process significantly more captions (56 and 65, respectively), indicating that LVNet achieves a superior balance between efficiency and accuracy.
-We present extended zero-shot evaluation results on IntentQA in Table A.8, comparing LVNet with prior zero-shot models across different reasoning categories: Why?, How?, and B.A.
+We present extended zero-shot evaluation results on IntentQA in Table˜A.13, comparing LVNet with prior zero-shot models across different reasoning categories: Why?, How?, and B.A.
 (Before/After).
@@
 Similarly, LangRepo and LLoVi, which process 90 captions, achieve overall scores of 59.1% and 64.0%, respectively, further demonstrating LVNet’s caption efficiency.
-To ensure fairness, models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in grey or downplayed in light green in Table A.7 and Table A.8.
-We adopt the reported GPT-4o results for VideoAgent and VideoTree in Section 4 and LABEL:tab:ablation:numberofframecaptions from the VCA Yang et al.
-(2024), but do not compare directly against VCA for two reasons: (1) the VCA paper does not provide code or implementation details (e.g., inference speed), making replication infeasible; and (2) its reported results cover only a subset of Egoschema, preventing a fair comparison to our approach on the full-scale EgoSchema, NExT-QA, and IntentQA.
-Overeall, these clarifications, alongside the results in Table A.7 and Table A.8, underscore LVNet’s effectiveness in achieving high accuracy while maintaining computational efficiency.
+To ensure fairness, models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in grey or downplayed in light green in Table˜A.12 and Table˜A.13.
 Model Cap. Cau. (%) Tem. (%) Des. (%) All (%)
@@
 Tarsier (Wang et al., 2024a) - - - - 79.2
-VideoAgent (Wang et al., 2024b) 8.2 72.7 64.5 81.1 71.3
+VideoAgent (Wang et al., 2024d) 8.2 72.7 64.5 81.1 71.3
 MVU (Ranasinghe et al., 2024) 16 55.4 48.1 64.1 55.2
@@
 VFC (Momeni et al., 2023) 32 45.4 51.6 64.1 51.5
-SeViLA (Yu et al., 2024) 32 61.3 61.5 75.6 63.6
-VideoTree (Wang et al., 2024e) (56) 75.2 67.0 81.3 73.5
+SeViLA (Yu et al., 2024a) 32 61.3 61.5 75.6 63.6
+VideoTree (Wang et al., 2024g) (56) 75.2 67.0 81.3 73.5
 ProViQ (Choudhury et al., 2023) 60 - - - 64.6
@@
 IG-VLM (Kim et al., 2024) - - - - 65.3
-SeViLA (Yu et al., 2024) 32 - - - 60.9
-VideoTree (Wang et al., 2024e) (56) - - - 66.9
+SeViLA (Yu et al., 2024a) 32 - - - 60.9
+VideoTree (Wang et al., 2024g) (56) - - - 66.9
 LangRepo (Kahatapitiya et al., 2024) 90 62.8 62.4 47.8 59.1
@@
 LVNet (ours) 12 75.0 74.4 62.1 71.7
-Table A.8: Extended results on IntentQA (Li et al., 2023a).
+Table A.13: Extended results on IntentQA (Li et al., 2023).
 We compare LVNet against prior zero-shot models across different reasoning categories: Why?, How?, and B.A.
@@
 Numbers in parentheses () indicate the maximum number of frames used.
+Appendix A.3 Extended results on VideoMME
+[…]
+Qwen-VL (Bai et al., 2023) 7B/OS no no 4 N/A 37.8
+Qwen-VL 7B/OS no no 1800 N/A OOM
+LongVILA (Chen et al., 2024a) 8B/OS no no 128 N/A 38.8
+LongVILA 8B/OS no no 256 N/A 39.7
+LLaVA-OneVision (Li et al., 2024) 7B/OS no no 8 N/A 43.8
+VideoChat-T (Zeng et al., 2024) 7B/OS no no 128 N/A 87.5 $\mathcal{O}(n)$ 41.9
+Frame-Voyager (Yu et al., 2024b) 7B/OS no no 128 N/A 93.7 $\mathcal{O}{(n\textsuperscript{r})}$ 48.9
+LLaVA-NexT-Video (Zhang et al., 2024a) 34B/OS no no 32 N/A 44.3
+Frame-Voyager 34B/OS no no 128 N/A 93.7 $\mathcal{O}{(n\textsuperscript{r})}$ 51.2
+InternVL2 (Chen et al., 2024c) 34B/OS no no 16 N/A 52.6
+LVNet (DeepSeek-v3) 37B/OS yes yes 1800 24 $\mathcal{O}(n)$ 53.1
+VideoAgent+GPT4o <1.8T/$P$ yes yes 24 $\mathcal{O}(m)$ 51.3
+VideoTree+GPT4o <1.8T/$P$ yes yes 300 24 92.0 $\mathcal{O}{(n\textsuperscript{r})}$ 52.3
+GPT4o direct <1.8T/$P$ yes yes 1800 1800 API Error
+LVNet (GPT-4o) <1.8T/$P$ yes yes 1800 24 98.7 $\mathcal{O}(n)$ 53.9
+LLaVA-Video (Zhang et al., 2024b) 72B/OS no no 64 N/A 61.5
+Qwen2-VL (Wang et al., 2024c) 72B/OS no no 768 N/A 62.2
+InternVL2.5 (Chen et al., 2024b) 72B/OS no no 64 N/A 62.6
+Table A.14: Detailed Comparison on VideoMME (Fu et al., 2024) .
+This table expands on Table˜3 by including two-stage (TS), video-level training free (VT Free), number of input frames used, two-stage captions fed to LLM (TS Cap.), ratio of frames/captions provided to the LLM relative to input frames (FR), and frame selection complexity.
+“OS” denotes open-source, "OOM” stands for out-of-memory.
+$\mathcal{O}(m)$ denotes the complexity proportional to the number LLM calls to predict keyframe timelines.
+The bottom 3 results are from benchmark leaderboard; we are unable to replicate these on our compute resources.
+Our LVNet variants require no video-level training yet achieve competitive results on very long videos.
+Table˜A.14 provides an extended comparison of both single-stage and two-stage approaches on the VideoMME long-video benchmark.
+In contrast to single-stage methods, which typically require costly video-level training on large datasets which has the risk of containing evaluation datasets.
+LVNet is video-level training free and thus retains broader generality.
+Critically, single-stage pipelines often become infeasible for long videos because they must process entire sequences—potentially exceeding 1,000 frames—through a heavy vision-language model (VLM) or LLM.
+Keyframe-selection approaches mitigate this challenge by filtering a minimal set of relevant frames before any large-model processing.
+As shown in Table˜A.14, LVNet can handle up to 1,800 frames by reducing them to just a few dozen keyframes, thereby remaining both memory- and compute-efficient while still achieving strong accuracy.
+For Qwen-VL and GPT4o, we demonstrate an ablation with 1,800 frames to highlight how easily single-stage methods can run out of memory (OOM) under 4 NVIDIA RTX A5000 GPUs or encounter GPT API errors.
+Moreover, our approach achieves strong performance with both open-source and proprietary LLM backbones, all without requiring any video-level training.
+LVNet+DeepSeek-V3 outperforms these single-stage keyframe selection models (VideoChat-T, Frame-Voyager) with equal or less frame selection complexity and outperforms single-stage VideoLLMs with a similarly sized open-source LLM.
+In a direct two-stage head-to-head setting that uses GPT-4o for both frame captioning and LLM reasoning and enforces the same 24-caption budget, LVNet further surpasses VideoAgent and VideoTree by +2.6% and +1.6%, respectively.
+Taken together, these results show LVNet’s effectiveness over both single-stage and two-stage video methods.
+The accuracy comes from each model’s recommended configuration and matched budgets for fair comparison.
+Appendix A.4 Qualitative Analysis of Hierarchical Keyframe Selector
+Figure A.6: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively.
+LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer.
+The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
+We compare open-ended responses of LVNet and the uniform sampling method in Figure˜A.6 to understand the effectiveness of the hierarchical keyframe selector in LVNet.
+The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively.
+LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames.
+When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, measuring tape), which is the correct answer, leading LVNet to choose the correct option.
+In contrast, the uniform sampling method selects frames at 0, 16, and 32 seconds and generates captions based on those frames.
+Similarly, when we feed the concatenated captions to the LLM to answer the same question, the output identifies only one activity—welding tools—resulting in the selection of the incorrect option.
+This example highlights the importance of keyframe selection and demonstrates the effectiveness of hierarchical keyframe selection in LVNet.
 Appendix A.5 Algorithms in Detail
-Our algorithms are presented in full detail in Algorithm 1, Algorithm 2, and Algorithm 3.
-TSC in Algorithm 1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets.
+Our algorithms are presented in full detail in Algorithm˜1, Algorithm˜2, and Algorithm˜3.
+TSC in Algorithm˜1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets.
 Within each of the $n$ sets, we uniformly sample $\leq\tau$ frames, obtaining a total of $T_{a}\leq\tau\times n$ frames.
 For example, LVNet sets $\psi=5,\lambda=12,\tau=18$, resulting in approximately $n\sim 25$ and $T_{a}\sim 390$ on the EgoSchema dataset.
-CKD in Algorithm 2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16.
+CKD in Algorithm˜2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16.
 LVNet employs $L=32,len(K)\leq 25$ on the EgoSchema dataset.
-FKD in Algorithm 3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally.
+FKD in Algorithm˜3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally.
 It groups frames sequentially into visual templates, each consisting of $N$ frames.
@@
 2:for all $img^{i}$ in $\textbf{List}_{frame}$ do
-43:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
-4:   $\textbf{List}_{feat}$.insert($\mathbf{F}^{i}$)
+[…]
+[…]
 55:end for
 6:for all $\mathbf{F}^{i}$ in $\textbf{List}_{feat}$ do
-67:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
-8:   $\textbf{M}_{dist}$.insert($\textbf{List}_{dist}$)
+[…]
+[…]
 79:end for
 10:while length of $\textbf{List}_{index}>\psi$ do
-11:   $\textbf{List}_{sample}\leftarrow\emptyset$
-12:   $\textbf{List}_{\delta}\leftarrow\emptyset$
-813:   $i\leftarrow\textbf{List}_{index}$.pop(00)
-914:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
-1015:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
-1116:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 17:for all prob in $\textbf{p}^{i}$ do
 18:  if prob < $\beta$ then
-19:         $\textbf{List}_{selected}$.insert($\textbf{index}(\texttt{prob},\textbf{p}^{i})$)
+[…]
 20:  end if
@@
 22:for all $\gamma$ in $\textbf{List}_{selected}$ do
-1323:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
+1323:  $\delta$ $\leftarrow$ $\gamma~th$ value in $\textbf{List}_{index}$
 24:  $\textbf{List}_{\delta}$.insert($\delta$)
@@
 1426:end for
-1527:   $\textbf{List}_{\delta}$.insert($i$)
-1628:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
-1729:   $\textbf{sort}(\textbf{List}_{sample})$
+[…]
+[…]
+[…]
 30:for all $frame^{j}$ in $\textbf{List}_{frame}$ do
@@
 3
-42:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
-53:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
+[…]
+[…]
 4:Initialize $\mathbf{P}_{best}$ as an empty list
@@
 1
+21:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, similarity score list $\mathbf{S}$, total length $L$, number of low similarity indices $K$, number of frames per visual template NN, number of keyframes selected per visual template $M$, function to sort by similarity $\textbf{sort}(\mathbf{S})$, function to order indices temporally $\textbf{temporal\_order}()$
+3
 […]
-3
-42:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
 […]
-64:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
 […]
-86:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
+[…]
+[…]
 […]
 8:for each $\mathbf{set}\in\mathbf{sets}$ do
-9:   $\mathbf{I}_{selected}\leftarrow\texttt{select\_top\_M}(\mathbf{set},M)$
+[…]
 1010:end for
-$1$:return $\mathbf{I}_{selected}$
+1111:return $\mathbf{I}_{selected}$
 Algorithm 3 Fine Keyframe Detection Process (FKD)
@@
 Each template image contains 8 images, and their order is described in language (e.g.
-top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure A.7.
+top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure˜A.7.
 Figure A.7: Prompt for Fine Keyframe Detection: The figure illustrates the input image, the prompt provided to the VLM, and the output.
@@
 The output displays the top three selected frames and their associated keywords.
+Appendix A.7 Integration to Existing Methods
+Our LVNet has been successfully integrated with other works for evaluation on long video benchmarks.
+For example, in Ranasinghe et al.
+(2024) LVNet is integrated with their proposed MVU to gain further performance boosts on the EgoSchema and NextQA benchmarks.
 Appendix A.8 Comparison with Other Keyframe Selection Methods
@@
 Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.
-Figure A.5 visualizes the differences of the keyframe selection mechanism bewtween LVNet and VideoAgent.
+Figure˜5 visualizes the differences of the keyframe selection mechanism bewtween LVNet and VideoAgent.
 On the left, LVNet begins with uniformly sampled frames and filters them through multiple stages, resulting in a non-uniform distribution of frames over time.
-First, the temporal scene clustering (TSC) selects somes frames that represent temporally distinct activities.
+First, the temporal scene clustering (TSC) selects some frames that represent temporally distinct activities.
 Next, the coarse keyframe detector (CKD) targets frames most relevant to the question.